### v1

Abstract
Distributions in spatial model often exhibit localized features. Intuitively, this locality implies a low intrinsic dimensionality, which can be exploited for efficient approximation and computation of complex distributions. However, existing approximation theory mainly considers the joint distributions, which does not guarantee that the marginal errors are small. In this work, we establish a dimension independent error bound for the marginals of approximate distributions. This $\ell_{\infty}$-approximation error is obtained using Stein’s method, and we propose a $\delta$-locality condition that quantifies the degree of localization in a distribution. We also show how $\delta$-locality can be derived from different conditions that characterize the distribution’s locality. Our $\ell_{\infty}$ bound motivates the localization of existing approximation methods to respect the locality. As examples, we show how to use localized likelihood-informed subspace method and localized score matching, which not only avoid dimension dependence in the approximation error, but also significantly reduce the computational cost due to the local and parallel implementation based on the localized structure.
1 Introduction
Approximation of distributions is a fundamental problem in applied mathematics and computer science. With the development of modern machine learning methods [15, 9, 21, 11], significant advancements are achieved in approximation of complex distributions. Compared to vast empirical success, concrete mathematical understanding of these methods is still lacking, especially in the context of high-dimensionality.
In this paper, we develop a $\ell_{\infty}$-approximation theory for a large class of high-dimensional distributions that exhibit localized features, explaining how such distributions can be approximated efficiently without suffering from the curse of dimensionality. These localized distributions are common in spatial models and graphical models arising from many statistical physics and computer science problems. A typical example is the Ising model where there are only local interactions.
Intuitively, such localized structure implies that the distribution is intrinsically low-dimensional, which can be exploited for efficient approximation and computation. We formalize this intuition and proves a dimension independent error bound for the marginals of approximate distributions. More precisely, we show in 2.4 that under some locality assumption, the $1$-Wasserstein marginal approximation error is bounded by local perturbations of the log density:
$\max_{i}W_{1}(\pi_{i},\pi_{i}^{\prime})\leq\delta\cdot\max_{i}\int\left|\nabla_{i}\log\pi(x)-\nabla_{i}\log\pi^{\prime}(x)\right|\pi(x)\mathrm{d}x.$
Note that the choice of the log density is motivated by its foundational role in many state-of-the-art approximation and sampling methods, such as the score matching [14], variational autoencoder [15] and diffusion models [21, 11]. Approximation error in terms of the log density is well documented in the literature, but most only consider the joint distribution, which in general is dimension dependent. One typical example is the Otto-Villani inequality [19]:
$W_{2}^{2}(\pi,\pi^{\prime})\leq c^{-2}\int\left|\nabla\log\pi(x)-\nabla\log\pi^{\prime}(x)\right|^{2}\pi(x)\mathrm{d}x.$
These bounds does not guarantee that the marginal error is small, and becomes less informative when one is only interested in the statistics of the marginals. Our $\ell_{\infty}$-approximation result fills the gap by exploiting the locality of the distribution and avoid the dimension dependence here.
Locality in the distributions is well-known and exploited in the communities like spatial statistics [25, 6, 20], data assimilation [13, 10] and image processing [27]. The key observation is that the correlation of parameters at large distances is neglectable in many spatial models. Therefore, modeling and computing such distributions can be done in an approximate manner, by artificially removing long-range correlations, thus greatly improving the computational efficiency. Such technique is called Vecchia approximation [25] in spatial statistics, and localization [13] in data assimilation. Despite the extensive application studies, the mathematical characterization of this approximation is rarely explored. With this in mind, we discuss in Section 2.3 the problem of approximating localized distributions by undirected graphical models (also known as the Markov random fields [16, 27]), i.e. one aims to find local potentials $u_{i}$ that depends only the neighboring sites $N_{i}$ of the site $i$, such that
$\pi(x)\approx\frac{1}{Z}\exp\left(\sum_{i}u_{i}(x_{N_{i}})\right).$
Note this is the prototype math formulation of the Vecchia approximation.
In the sampling literature, there are also recent work [17, 23, 8, 2] exploiting the locality for efficient sampling. [17] proposed to apply the localization technique in MCMC sampling to solve Bayesian inverse problems and obtained a dimension independent convergence rate. [23] extended this idea and proved a dimension independent convergence rate of the Metropolis-Adjusted-Langevin-Algorithm-within-Gibbs (MLwG) for distributions with a sparse condition structure (equivalent to a sparse graphical model). [8] made a initial trial to obtain $\ell_{\infty}$-approximation error for a image deblurring problem, based on which we further develop the theory to the general case in this paper. We also notice a concurrent work [2] that shows the marginals of unadjusted Langevin algorithm converges to the target at a convergence rate with logarithm dependence on the dimension. They introduced a $\ell_{\infty}$-type Wasserstein distance, which follows a similar focus on the marginals as in our work.
In this paper, we propose the concept of $\delta$-locality in 2.1 to quantifies how a distribution is localized. We will use the $\delta$-locality to obtain $\ell_{\infty}$-approximation error by applying the Stein’s method [22]. We show how to obtain $\delta$-locality given different characterization of the localized distribution, including the sparse graphical model. The main technical contribution is to show that $\delta$ is dimension independent for typical localized distributions. The $\ell_{\infty}$-approximation bound motivates to localize existing approximation methods to respect the locality. As examples, we show how to localize the likelihood-informed subspace (LIS) method [4, 3, 5] and the score matching method [14, 21, 11]. Such localized approximation method not only yields a dimension independent $\ell_{\infty}$ error bound, but also greatly reduces the computational cost due to the local and parallel implementation based on the localized structure.
The paper is organized as follows. The main results are presented in Section 2, where we define the $\delta$-locality, proves the main approximation theory, and then discuss some useful extensions. The following three sections discuss three applications of the theory: the localized LIS method in Section 3, the localized score matching in Section 4, and the localized neural network surrogate in Section 5. In Section 6, we provide the main steps of the technical results for the $\delta$-locality.
Notations
  • Norms. For $v\in\mathbb{R}^{d}$, briefly denote the vector $2$-norm as $\left|v\right|$. Likewisely, for a matrix $A\in\mathbb{R}^{m\times n}$, $\left\|A\right\|=\sup_{v\neq 0}\frac{\left|Av\right|}{\left|v\right|}$ denotes the matrix $2$-norm. $\left\|A\right\|_{\rm F}$ denotes the Frobenius norm $\left\|A\right\|_{\rm F}^{2}=\sum_{i,j}A_{ij}^{2}$.
    For a function $f:\mathbb{R}^{d}\to\mathbb{R}$, denote $\left\|f\right\|_{L^{p}(\pi)}$ as the weighted $L^{p}$ norm $\left\|f\right\|_{L^{p}(\pi)}^{p}=\int_{\mathbb{R}^{d}}|f(x)|^{p}\pi(x)\mathrm{d}x$. For a vector-valued function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$, still denote $\left\|f\right\|_{L^{p}(\pi)}^{p}=\int_{\mathbb{R}^{d}}\left|f(x)\right|^{p}\pi(x)\mathrm{d}x$ and note here $\left|f(x)\right|$ is the vector $2$-norm without specification.
  • Matrix. Denote $A\succeq B$ if $A-B$ is positive semi-definite, and $A\preceq B$ if $B\succeq A$.
  • Probability. Denote $\mathcal{P}_{1}(\mathbb{R}^{d})$ as the set of probability distributions in $\mathbb{R}^{d}$ with finite first moment. The $1$-Wasserstein distance between two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is defined as
    $W_{1}(\pi,\pi^{\prime})=\inf_{\gamma\in\Pi(\pi,\pi^{\prime})}\int_{\mathbb{R}^{d}}\left|x-y\right|\mathrm{d}\gamma(x,y),$
    where $\Pi(\pi,\pi^{\prime})$ is the set of couplings of $\pi$ and $\pi^{\prime}$, i.e. if $(X,Y)\sim\gamma\in\Pi(\pi,\pi^{\prime})$, then $X\sim\pi,Y\sim\pi^{\prime}$.
    Denote conditional independence $X\perp\!\!\!\perp Y~{}|~{}Z$ if $p(x,y|z)=p(x|z)p(y|z)$.
  • Block decomposition. For the parameter $x\in\mathbb{R}^{d}$, we assume there is a fixed block decomposition
    $x=(x_{1},\dots,x_{b}),\quad\forall i\in[b],~{}x_{i}\in\mathbb{R}^{d_{i}},\quad d=\sum_{i=1}^{b}d_{i},$
    where we denote $[b]:=\{1,\dots,b\}$. Denote $-i:=[b]\backslash\{i\}$. For a distribution $\pi$, denote $\pi_{i}$ as the marginal distribution on $x_{i}$, i.e. $\pi_{i}(x_{i})=\int\pi(x_{i},x_{-i})\mathrm{d}x_{-i}$. Denote the partial derivatives $\nabla_{i}=\nabla_{x_{i}},\nabla_{ij}^{2}=\nabla_{x_{i}x_{j}}^{2}$.
2 $\ell_{\infty}$-approximation theory
2.1 $\delta$-locality
We begin by defining the $\delta$-locality of a distribution, which quantifies the locality of the distribution.
Definition 2.1.
A distribution $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is called $\delta$-localized for some constant $\delta>0$, independent of $d$, if for any $i\in[b]$ and $1$-Lipschitz function $\phi:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}$ such that $\int\phi(x_{i})\pi_{i}(x_{i})\mathrm{d}x_{i}=0$, the solution $u(x)$ to the Stein equation
$\Delta u(x)+\nabla\log\pi(x)\cdot\nabla u(x)=\phi(x_{i})$ (2.1)
satisfies the gradient estimate
$\left\|\nabla u\right\|_{\infty,1}:=\sum_{j=1}^{b}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq\delta.$ (2.2)
2.1 is a technical description of the localized structure of the distribution motivated by the Stein’s method [22]. We will show in 2.4 that the $\delta$-locality condition directly implies that local perturbation of the distribution only introduces dimension independent marginal error. We will see from the proof there that
$\nabla_{j}u(x)$ quantifies how modifications of $\nabla_{j}\log\pi(x)$ affect the marginal $\pi_{i}$.
We keep the technical definition here for the sake of generality. To make it more concrete, we introduce below two common assumptions that imply the $\delta$-locality.
Proposition 2.2.
$\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is $\delta$-localized if it satisfies any of the following conditions:
  1. (1)
    ($\ell_{1}$-condition) $H(x)=-\nabla^{2}\log\pi(x)$ is $c$-unifomly diagonal block dominant, i.e. there exists a matrix $M\in\mathbb{R}_{+}^{b\times b}$ s.t. $\forall i,j\in[b],i\neq j$,
    $H_{ii}(x)\succeq M_{ii}I_{d_{i}},\quad\|H_{ij}(x)\|\leq M_{ij};\quad\sum_{j\neq i}M_{ij}+c\leq M_{ii},$
    where $H_{ij}(x)$ denotes the $(i,j)$-th subblock of $H(x)$. Here $\delta$ can be taken as $\delta=c^{-1}$.
  2. (2)
    ($\ell_{2}$-condition) $H(x)=-\nabla^{2}\log\pi(x)$ has uniform block bandwidth $B>0$, and uniform lower and upper bounded $m$ and $M$ where $0<m\leq M<\infty$, i.e.
    $\forall|i-j|>B,\quad H_{ij}(x)=0;\quad mI\preceq H(x)\preceq MI,$
    Here $\delta$ can be taken as $\delta=\frac{2BM}{m^{2}}$.
Proofs can be found in Section 6.1.
Remarks 2.3.
(1) Note that the two conditions above both require that $\pi$ is log-concave. To see that the $\ell_{1}$-condition implies log-concavity, denote $M_{ij}^{\prime}:=2M_{ii}\mathbf{1}_{i=j}-M_{ij}$, then $M^{\prime}$ is $c$-diagonal dominant. Gerŝgorin discs theorem [12] then implies the smallest eigenvalue of $M^{\prime}$ is lower bounded by $c$, and thus
$u^{\rm T}H(x)u=~{}$ $\sum_{i,j}u_{i}^{\rm T}H_{ij}(x)u_{j}\geq\sum_{i}M_{ii}\left|u_{i}\right|^{2}-\sum_{i\neq j}M_{ij}\left|u_{i}\right|\left|u_{j}\right|$
[…]
(2) The two conditions characterize the locality of the distribution in different ways. The $\ell_{1}$-condition uses diagonal dominance of the Hessian, which implies that any block is mostly influenced by itself. The $\ell_{2}$-condition directly uses bandedness of Hessian to characterize the localized structure.
2.2 $\ell_{\infty}$-approximation error
Theorem 2.4.
Consider two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Assume $\pi^{\prime}$ is $\delta$-localized, then the marginal $W_{1}$-distance of $\pi,\pi^{\prime}$ satisfies
$\max_{i\in[b]}W_{1}(\pi_{i},\pi_{i}^{\prime})\leq\delta\cdot\max_{j\in[b]}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)},$ (2.3)
where $\pi_{i}$ and $\pi_{i}^{\prime}$ are the marginals of $\pi$ and $\pi^{\prime}$ on $x_{i}$ respectively.
Proof.
By Kantorovich duality [26],
$W_{1}(\pi_{i},\pi_{i}^{\prime})=\sup_{\phi_{0}\in{\rm Lip}_{1}}\int\phi_{0}(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i},$
where ${\rm Lip}_{1}$ denotes the $1$-Lipchitz function class. Fix any $\phi_{0}\in{\rm Lip}_{1}$. Denote
$\phi(x_{i})=\phi_{0}(x_{i})-\int\phi_{0}(x_{i})\pi^{\prime}(x_{i})\mathrm{d}x_{i}.$
Then
$\int\phi_{0}(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}=~{}$ $\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}$
[…]
Here we use $\pi(x_{i})=\int\pi(x_{i},x_{-i})\mathrm{d}x_{-i}$. Let $u(x)$ solve the Stein equation
$\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)=\phi(x_{i}).$
Now by integration by parts,
$\int\phi(x_{i})\pi(x)\mathrm{d}x=~{}$ $\int\left(\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)\right)\pi(x)\mathrm{d}x$
[…]
By assumption, the solution satisfies the gradient estimate (2.2). Therefore,
$\int\phi(x_{i})\pi(x)\mathrm{d}x\leq~{}$ $\sum_{j=1}^{b}\left\|\nabla_{j}u\right\|_{L^{\infty}}\int\left|\nabla_{j}\log\pi^{\prime}(x)-\nabla_{j}\log\pi(x)\right|\pi(x)\mathrm{d}x$
[…]
Therefore, we obtain
$\max_{i}W_{1}(\pi_{i},\pi_{i}^{\prime})\leq~{}$ $\max_{i}\max_{\phi_{0}\in{\rm Lip}_{1}}\int\phi_{0}(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}$
[…]
This completes the proof. ∎
2.3 Approximation by sparse graphical models
In this section, we consider a special case of approximating a distribution $\pi$ by an undirected graphical model [16, 27]. The undirected graphical model, also called Markov random field, models the local dependence of the distribution by an undirected graph. It requires that any nonadjacent two components $X_{i},X_{j}$ are conditionally independent given the rest of the components.
Consider an undirected graph $G=(V,E)$ where $|V|=b$. We attach each vertex in $V$ with a block index $i$ uniquely. Denote $N_{j}=\{k\in V:\overline{jk}\in E\}$ as the neighboring vertices of $j$. Here we require $E$ includes all the self-loop in $G$, i.e. $\forall j\in G,\overline{jj}\in E$.
We say a distribution $\pi(x)\in\mathcal{P}(\mathbb{R}^{d})$ has dependence graph $G$ if
$\forall j,k\in V,~{}k\notin N_{j}~{}\Rightarrow~{}x_{j}\perp\!\!\!\perp x_{k}~{}|~{}x_{N_{j}^{-}},$ (2.4)
[…]
$k\notin N_{j}~{}\Rightarrow~{}\nabla_{jk}^{2}\log\pi(x)=0.$
Note it generalizes the banded condition on the Hessian in 2.2 by using the sparse graphical structure. Parallelly, one can show that under certain sparsity condition on $G$, a distribution $\pi$ with dependence graph $G$ is $\delta$-localized.
To specify the sparsity condition, we denote the $q$-th neighbor $N_{j}^{q}~{}(q\in\mathbb{Z}_{+})$ as
$N_{j}^{q}:=\{k\in V:d_{G}(j,k)\leq q\},$
where $d_{G}(j,k)$ is the shortest path distance between $j$ and $k$ in $G$, i.e. the minimum number of edges that must be traversed to go from $j$ to $k$. For convenience, we take $d_{G}(j,j)=0$, and $d_{G}(j,k)=\infty$ if $j,k$ are not connected.
Definition 2.5.
A $s$-sparse graph $G$ with ambient dimension $r\in\mathbb{Z}_{+}$ is such that
$\forall j\in V,\quad|N_{j}^{q}|\leq 1+sq^{r}.$
We also call $G$ a $s$-sparse graph without specifying the ambient dimension.
Remark 2.6.
The definition of sparse graph is different from that in graph theory. Roughly speaking, the sparsity $s$ denotes the size of the immediate neighbor, and the dimension $r$ controls the growth rate of the neighbor size with respect to the distance. Note we require it grows at most polynomially to ensure the locality. A typical example of such sparse graph is the lattice model $\mathbb{Z}^{r}$. A naive bound for $\mathbb{Z}^{r}$ is $|N_{j}^{q}|\leq(2q+1)^{r}<1+(3q)^{r}$. So that $\mathbb{Z}^{r}$ is $3^{r}$-sparse with ambient dimension $r$.
Proposition 2.7.
Let $G=(V,E)$ be a $s$-sparse graph with ambient dimension $r$. Suppose $\pi\in\mathcal{P}(\mathbb{R}^{d})$ is a distribution with dependence graph $G$, and satisfies
$\forall x\in\mathbb{R}^{d},\quad mI\preceq-\nabla^{2}\log\pi(x)\preceq MI.$
Then $\pi$ is $\delta$-localized with $\delta=\frac{sr!\kappa^{r}}{m}$ where $\kappa=\frac{M}{m}$. As a consequence, for any $\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$,
$\max_{j\in V}W_{1}(\pi_{j},\pi_{j}^{\prime})\leq\frac{sr!\kappa^{r}}{m}\max_{k\in V}\left\|\nabla_{k}\log\pi-\nabla_{k}\log\pi^{\prime}\right\|_{L^{1}(\pi^{\prime})},$
Proofs can be found in Section 6.2. Note this recovers the second result in 2.2, where the bandedness condition corresponds to a $2B$-sparse graph with ambient dimension $r=1$.
2.4 Marginal error of multiple blocks
In 2.4, we show that the marginal distance of $\pi^{\prime}$ and a $\delta$-localized distribution $\pi$ on one block $x_{i}$ is dimension independent. Here we extend this result to the case of multiple blocks. This result provides further control on the correlation between different blocks in $\pi^{\prime}$, which may not be captured by 2.4.
Theorem 2.8.
Consider two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Assume $\pi^{\prime}$ satisfies any one of the conditions in 2.2 or 2.7, then for any index set $I\subset[b]$, the marginal $W_{1}$-distance of $\pi,\pi^{\prime}$ on $x_{I}:=(x_{i})_{i\in I}$ satisfies
$W_{1}(\pi_{I},\pi_{I}^{\prime})\leq\delta|I|\cdot\max_{j\in[b]}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.$ (2.5)
Here $\delta$ can be taken as the same constant in 2.2 or 2.7.
Proofs can be found in Section 6.3. The implications of the theorem is two-fold. First, for blocks $i,j$ at large distance, the correlation of $x_{i},x_{j}$ under $\pi$ is neglectable, and $\pi_{i,j}$ are almost independent: $\pi_{i,j}\approx\pi_{i}\pi_{j}$. The above theorem says that $\pi^{\prime}_{i,j}$ is close to $\pi_{i,j}$, so that $x_{i},x_{j}$ are also almost independent under $\pi^{\prime}$. Conversely, for neighboring blocks $i,j$, $\pi^{\prime}_{i,j}$ can capture the correlation between $x_{i},x_{j}$ in $\pi$.
3 Application I: localized LIS method
Likelihood-informed subspace (LIS) method [4, 3, 5] is a dimension reduction technique to sample efficiently from a high-dimensional distribution of the form
$\pi(x)=\frac{1}{Z}l(x)\pi_{0}(x),$ (3.1)
where $\pi_{0}(x)$ is the prior distribution, and $l(x)$ is the likelihood function. Sampling from such distributions is a central problem in Bayesian inference. The key in the LIS method is the low rank structure of the likelihood function $l(x)$, making it possible to be approximated by a rigde function
$l(x)\approx l_{r}(x_{r}),\quad x_{r}=P_{r}x,$
where $P_{r}$ is a projection to a low-dimensional subspace $\mathcal{X}_{r}\subset\mathcal{X}$. The LIS method propose to approximate the target distribution by
$\pi(x)\approx\pi_{r}(x)\propto l_{r}(x_{r})\pi_{0}(x),$
which, with properly chosen informed subspace $\mathcal{X}_{r}$ and ridge approximation of $l(x)$, provides certified approximation error bounds [5]. With such low-rank approximation, the sampling complexity is greatly reduced, and the method has been successfully applied in many high-dimensional problems [4, 3].
3.1 Localized LIS method
In LIS, the informed subspace is constructed globally, which may not capture the spatial structure in the target distribution. For distributions with sparse dependence graph, a locally defined informed subspace can be a better choice. Based on this observation, we propose to use localized likelihood-informed subspace (LLIS), which incorporates domain decomposition in LIS method. In essence, we decompose the parameters into small blocks, and apply LIS in each block to construct the LLIS.
In the following, we show how to apply LLIS when the prior is a Gaussian distribution, and establish a dimension independent error for LLIS approximation under the sparse graphical assumption.
LLIS construction
We propose to apply the LLIS approximation under a transformed representation so that the prior is a standard Gaussian, i.e. consider the change of variables
$\tilde{x}=C^{-1/2}x.$
Then $\tilde{x}\sim\tilde{\pi}_{0}=\mathcal{N}(0,I_{d})$ if $x\sim\pi_{0}$. Under the transformed representation $\tilde{x}$, we aim to find the LLIS of the product form
$\widetilde{\mathcal{X}}_{r}=\bigoplus_{j\in V}\widetilde{\mathcal{X}}_{r,j},$
where $\widetilde{\mathcal{X}}_{r,j}\subset\mathbb{R}^{d_{j}}$ denotes the subspace of the reduced paramater in block $j$. To determine $\widetilde{\mathcal{X}}_{r,j}$, consider the following diagnostic matrix
~$\begin{split}\widetilde{\mathcal{M}}=~{}&\mathbb{E}_{\tilde{x}\sim\tilde{\pi}}\left[\tilde{\nabla}\log\tilde{l}(\tilde{x})(\tilde{\nabla}\log\tilde{l}(\tilde{x}))^{\rm T}+\tilde{\nabla}^{2}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}^{2}\log\tilde{l}(\tilde{x})\right]\\ =~{}&C^{1/2}\cdot\mathbb{E}_{x\sim\pi}\left[\nabla\log l(x)(\nabla\log l(x))^{\rm T}+\nabla^{2}\log l(x)\cdot C\cdot\nabla^{2}\log l(x)\right]\cdot C^{1/2},\end{split}$ (3.2)
where we denote the transformed quantities with a tilde, to be specific,
$\tilde{x}=C^{-1/2}x,\quad\tilde{\pi}(\tilde{x})=(\det C)^{-\frac{1}{2}}\pi(x),\quad\tilde{l}(\tilde{x})=l(x),\quad\tilde{\nabla}=\nabla_{\tilde{x}},\quad\tilde{\nabla}^{2}=\nabla^{2}_{\tilde{x}\tilde{x}}.$
Notice compared to the existing LIS methods [5] where only first order information is used, here we also incorporates the second order information in the diagnostic matrix.
With the diagnostic matrix, we define $\widetilde{\mathcal{X}}_{r,j}$ to be the rank-$r_{j}$ principle subspace of ~$\widetilde{\mathcal{M}}_{jj}\in\mathbb{R}^{d_{j}\times d_{j}}$, the diagonal block of $\widetilde{\mathcal{M}}$ for block $j$. To be specific,
$\widetilde{\mathcal{X}}_{r,j}=\text{span}\{u_{1},\dots,u_{r_{j}}\},$
where $u_{1},\dots,u_{r_{j}}$ are the eigenvectors of ~$\widetilde{\mathcal{M}}_{jj}$ with the largest $r_{j}$ eigenvalues. Here $r_{j}=\text{\rm rank}(\widetilde{\mathcal{X}}_{r,j})$ is the number of informed directions we choose. It can be determined, for instance, by (see [5])
$r_{j}=\min\left\{1\leq r\leq d_{j}:\sum_{k=1}^{r}\lambda_{k}(\widetilde{\mathcal{M}}_{jj})\geq(1-\epsilon)\text{\rm tr}\left[\widetilde{\mathcal{M}}_{jj}\right]\right\},$
where $\lambda_{k}(\widetilde{\mathcal{M}}_{jj})$ denotes the $k$-th largest eigenvalue of ~$\widetilde{\mathcal{M}}_{jj}$, and $\epsilon\in(0,1)$ is some prescribed threshold.
Given $\widetilde{\mathcal{X}}_{r}$, denote accordingly the complementary space
$\widetilde{\mathcal{X}}_{\bot,j}=(\widetilde{\mathcal{X}}_{r,j})^{\bot},\quad\widetilde{\mathcal{X}}_{\bot}=\bigoplus_{j\in V}\widetilde{\mathcal{X}}_{\bot,j}.$
Denote the reduced parameter $\tilde{x}_{r}$, its complement $\tilde{x}_{\bot}$ and their counterparts in the original representation as
$\tilde{x}_{r}=(\tilde{x}_{r,j})_{j\in V},\quad\tilde{x}_{r,j}=\tilde{P}_{r,j}\tilde{x}_{j},\quad x_{r}:=C^{1/2}\tilde{x}_{r}\in\mathcal{X}_{r}:=C^{1/2}\widetilde{\mathcal{X}}_{r}.$
$\tilde{x}_{\bot}=\tilde{x}-\tilde{x}_{r}=(\tilde{x}_{\bot,k})_{j\in V},\quad\tilde{x}_{\bot,k}=\tilde{x}_{j}-\tilde{x}_{r,j}=\tilde{P}_{\bot,k}\tilde{x}_{j},\quad x_{\bot}:=C^{1/2}\tilde{x}_{\bot}\in\mathcal{X}_{\bot}:=C^{1/2}\widetilde{\mathcal{X}}_{\bot}.$
Here $\tilde{P}_{r,j}$ is the orthogonal projector onto $\widetilde{\mathcal{X}}_{r,j}$, and $\tilde{P}_{\bot,k}$ is its orthogonal complement. Note $x_{r}$ and $x_{\bot}$ are not orthogonal in the usual sense, but it holds that
$x_{r}^{\rm T}C^{-1}x_{\bot}=\tilde{x}_{r}^{\rm T}\tilde{x}_{\bot}=0.$
That is, $x_{r}$ and $x_{\bot}$ are independent under the prior $\pi_{0}=\mathcal{N}(0,C)$.
LLIS approximation
As in the LIS method, we construct the approximate likelihood $l_{r}(x_{r})$ by averaging $l(x)$ in the logarithm regime over the uninformed subspace $\mathcal{X}_{\bot}$, and approximate the target distribution by
$\begin{split}&\pi_{r}(x)\propto l_{r}(x_{r})\pi_{0}(x),\quad x_{r}=P_{r}x,\\ \text{where}\quad&\log l_{r}(x_{r})=\int\log l(x_{r}+x_{\bot})\pi_{0}(x_{\bot})\mathrm{d}x_{\bot}.\end{split}$ (3.3)
Note $x_{\bot}$ and $x_{r}$ are independent under $\pi_{0}$, so that $\pi_{0}(x_{\bot}|x_{r})=\pi_{0}(x_{\bot})$. Also note in the transformed representation, the approximation takes the same form,
$\begin{split}\log\tilde{l}_{r}(\tilde{x}_{r}):=~{}&\log l_{r}(C^{1/2}\tilde{x}_{r})=\int\log l(C^{1/2}\tilde{x}_{r}+x_{\bot})\pi_{0}(x_{\bot})\mathrm{d}x_{\bot}\\ =~{}&\int\log\tilde{l}(\tilde{x}_{r}+C^{-1/2}x_{\bot})\pi_{0}(x_{\bot})\mathrm{d}x_{\bot}=\int\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}.\end{split}$ (3.4)
Here we choose to average in the logarithm regime for the sake of simplicity of the theoretical results. In practice, one can also consider other averaging schemes, see [5].
3.2 Approximation error
We show the marginal approximation error of the LLIS method is dimension independent under the sparse graphical assumption. To be specific, we assume
Assumption 3.1.
The target distribution is of the form (3.1), and
  • $\pi_{0}(x),\pi(x)$ have dependence graph $G$ that is $s$-sparse with ambient dimension $r$.
  • $\pi_{0}(x)=\mathcal{N}(0,C)$, and there exists $0<m\leq M<\infty$ s.t.
    $mI\preceq C^{-1}\preceq MI,\quad mI\preceq-\nabla^{2}\log\pi(x)\preceq MI.$
As in [4, 5], we will bound the approximation error of LLIS by the eigenvalue residues of the modified diagnostic matrices (3.2). For simplicity, denote the eigenvalue residue for a positive semidefinite matrix $C\in\mathbb{R}^{d\times d}$ and a subspace $V_{r}\subset\mathbb{R}^{d}$ as
$\mathcal{R}(C,V_{r}):=\text{\rm tr}\left[P_{\bot}CP_{\bot}\right],$ (3.5)
where $P_{\bot}$ is the orthogonal projector onto $V_{r}^{\bot}$.
Theorem 3.2.
Assume the target distribution $\pi$ satisfies 3.1. Consider its LLIS approximation $\pi_{r}$ (3.3). Then there exists a dimension independent constant $C_{\pi}$ s.t. the marginal error of the LLIS approximation satisfies
$\max_{j}W_{1}(\pi_{j},\pi_{r,j})\leq C_{\pi}\max_{k\in V}\left(\mathcal{R}(\widetilde{\mathcal{G}}_{k},\widetilde{\mathcal{X}}_{r,k})+\sum_{j\in V}\mathcal{R}(\widetilde{\mathcal{H}}_{j,k},\widetilde{\mathcal{X}}_{r,j})\right)^{1/2},$ (3.6)
where we denote $\mathcal{R}$ as the residue defined in (3.5), and the diagnostic matrices
$\begin{split}\widetilde{\mathcal{G}}_{k}:=~{}&\mathbb{E}_{\tilde{x}\sim\tilde{\pi}_{r}}\left[\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x}))^{\rm T}\right].\\ \widetilde{\mathcal{H}}_{j,k}:=~{}&\mathbb{E}_{\tilde{x}\sim\tilde{\pi}_{r}}\left[\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{k}\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})\right].\end{split}$ (3.7)
Proofs can be found in Appendix B.
Remarks 3.3.
(1) We would like to compare the right rand side of (3.6) with the diagnostic matrix (3.2):
  • The expectation in (3.7) is taken w.r.t. the approximate distribution $\tilde{\pi}_{r}$, while in (3.2) it is taken w.r.t. the target distribution $\tilde{\pi}$. So that 3.2 is a posteriori error estimate instead of a priori estimate.
  • (3.2) can be written as
    ~$\widetilde{\mathcal{M}}_{jj}=\mathbb{E}_{\tilde{x}\sim\tilde{\pi}}\left[\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x}))^{\rm T}+\sum_{k\in V}\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{k}\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})\right].$
    For the second term here, the summation is taken on the second index ($k$) in ~$\tilde{\nabla}_{j}\tilde{\nabla}_{k}$, while in (3.6), it is taken on the first index ($j$) in ~$\tilde{\nabla}_{j}\tilde{\nabla}_{k}$.
(2) We would also like to comment that the summation $\sum_{j\in V}$ in (3.6) is still a dimension independent term. Since the second order derivatives ~$\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})$ decay fast with $d_{G}(j,k)$, which is a result of the sparsity graphical assumption.
4 Application II: localized score matching
The score function of a distribution $p(x)$ is defined as $s(x):=\nabla\log p(x)$. In score matching [14, 21], one uses variational method to approximate $p(x)$ by fitting the score function. Given a hypothesis class $\mathscr{P}=\{p_{\theta}(x):\theta\in\Theta\}$ of distributions parameterized by $\theta$, the score matching aims to minimize the Fisher divergence over $\theta$:
$D_{\rm Fisher}(p_{\theta},p):=\mathbb{E}_{X\sim p}\left|s_{\theta}(X)-s(X)\right|^{2},\quad s_{\theta}(x):=\nabla\log p_{\theta}(x).$
Though the Fisher divergence is not directly computable, a simple integration by parts gives a tractable loss function
$J(\theta)=\mathbb{E}_{X\sim p}\left[2\text{\rm tr}\left(\nabla s_{\theta}(X)\right)+\left|s_{\theta}(X)\right|^{2}\right].$
Here $J(\theta)=D_{\rm Fisher}(p_{\theta},p)-\mathbb{E}_{X\sim p}\left|\nabla\log p(X)\right|^{2}$, so that minimizing $J(\theta)$ is equivalent to minimizing the Fisher divergence. In practice, $J(\theta)$ is approximated by its empirical version: given data $(X_{1},\dots,X_{N})$ drawn from $p$, the empirical loss is
$\hat{J}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[2\text{\rm tr}\left(\nabla s_{\theta}(X_{i})\right)+\left|s_{\theta}(X_{i})\right|^{2}\right].$
The hypothesis class $\mathscr{P}$ is constructed such that the gradient of the score function $s_{\theta}(x)$ is easy to compute and optimize over, so that an approximate distribution $\hat{p}_{\theta}$ can be obtained efficiently by minimizing $\hat{J}(\theta)$.
Despite its simple implementation, fitting a generic score function in high-dimensional space is still computationally expensive. We propose to reduce the complexity by leveraging the localized structure, i.e. the target distribution is or can be approximated by a sparse graphical model. The locality in the model largely reduces the complexity of the score function, based on which we propose the localized score matching method.
4.1 Localized score matching
Factorization of sparse graphical models
The key motivation of localized score matching is the existence of local factorization of the log density for the sprase graphical models. Consider a distribution $p(x)$ with dependence graph $G=(V,E)$. By Hammersley-Clifford theorem [16], one can factorize
$\log p(x)=\sum_{j\in V}u_{j}(x_{N_{j}}).$
Note $u_{j}$ is a function of $x_{N_{j}}\in\mathbb{R}^{d_{N_{j}}}$, where we denote the local dimension
$d_{N_{j}}:=\sum_{k\in N_{j}}d_{k}.$
Suppose $G$ is $s$-sparse in the sense of 2.5, then the maximal local dimension satisfies
$d_{\rm loc}:=\max_{j}d_{N_{j}}\leq(1+s)\max_{j}d_{j}.$ (4.1)
In a sparse graphical model, the local dimension $d_{\rm loc}\ll d$, so that the complexity of $\log p(x)$ would be greatly reduced. This motivates to fit the target using distributions from a hypothesis class $\mathscr{P}_{G}$ that captures the localized structure. Specifically, we consider the hypothesis class
$\mathscr{P}_{G}=\left\{p_{\theta}(x)=\exp\left(\sum_{j\in V}u_{\theta,j}(x_{N_{j}})\right):u_{\theta,j}\in\mathcal{U}_{\theta,j}\right\},$
where $\mathcal{U}_{\theta,j}$ is some class of functions in $\mathbb{R}^{d_{N_{j}}}$ to be determined. The distributions in $\mathscr{P}_{G}$ have dependence graph $G$ by definition. Notice the size of the hypothesis space, measured by its metric entropy [24], is $O(\text{poly}(d)\exp(Cd_{\rm loc}))$, thus avoiding the curse of dimensionality.
Here we assume that the $G$ is prefixed based on prior knowledge of the target distribution. One can also learn the dependence graph from data, but we do not consider it here.
Localized loss function
We define the local score function as $s_{j}(x):=\nabla_{j}\log p(x)$. To fit the local score function, it suffices to minimize the local loss
$J_{j}(\theta):=\mathbb{E}_{X\sim p}\left[2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(X)\right)+\left|s_{\theta,j}(X)\right|^{2}\right],$ (4.2)
where $s_{\theta,j}(x)=\nabla_{j}\log p_{\theta}(x)$. Note for $p_{\theta}\in\mathscr{P}_{G}$, the local score $s_{\theta,j}$ can be computed locally by
$s_{\theta,j}(x)=\sum_{k\in N_{j}}\nabla_{j}u_{\theta,k}(x_{N_{k}}),$
which only concerns a small group of the local scores. To fit the whole score function, one could consider to minimize $\max_{v\in V}J_{j}(\theta)$. However, it is no longer equivalent to minimizing
$\max_{j\in V}\mathbb{E}_{X\sim p}\left|s_{\theta,j}(X)-s_{j}(X)\right|^{2}=\max_{j\in V}\left(J_{j}(\theta)+C_{j}\right),\quad C_{j}:=\mathbb{E}_{X\sim p}|s_{j}(X)|^{2}.$
The reason is that $C_{j}$ variates with $j$ and directly minimizing $\max_{j}J_{j}(\theta)$ may ignores those $j$ with larger $C_{j}$. Nevertheless, one can relax the problem to be a saddle point problem:
$\min_{\theta\in\Theta}J_{\rm loc}(\theta):=\min_{\theta\in\Theta}\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(J_{j}(\theta)+\lambda_{j}\right),$ (4.3)
Here $\Theta$ denotes the parameter space of the hypothesis class. Though $C_{j}$ is not directly accessible, one can replace it by some accessible lower bound of it in implementation. One also needs to replace the loss by the empirical version
$\begin{split}&\hat{J}_{\rm loc}(\theta)=\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(\hat{J}_{j}(\theta)+\lambda_{j}\right),\\ &\hat{J}_{j}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(X_{i})\right)+\left|s_{\theta,j}(X_{i})\right|^{2}\right],\end{split}$ (4.4)
where $X_{1},\dots,X_{N}$ are i.i.d. samples from $p$.
4.2 Statistical error
In this section, we aim to quantify rate the convergence of the minimizer $\hat{\theta}_{*}$ of the empirical loss (4.4) to the minimizer $\theta_{*}$ of the population loss (4.3), i.e.
$\theta_{*}=\mathop{\arg\min}_{\theta\in\Theta}J_{\rm loc}(\theta),\quad\hat{\theta}_{*}=\mathop{\arg\min}_{\theta\in\Theta}\hat{J}_{\rm loc}(\theta)$
in terms of the number of samples $N$ and the full dimension $d$. Take the hypothesis class $\mathscr{P}$ larger to include the true distribution, i.e. $p\in\mathscr{P}$, so that $p_{\theta_{*}}=p$. Then such convergence rate quantifies the sample complexity to learn the target distribution, i.e. how many samples are needed to approximate the target distribution within a certain error.
For simplicity, we consider the target distribution $p(x)$ supported in the unit interval $\mathbb{I}^{d}=[0,1]^{d}$, and take the following hypothesis class
$\begin{split}\mathscr{P}=\Big{\{}p_{\theta}(x)=\exp&\left(\sum_{j\in V}u_{\theta,j}(x_{N_{j}})\right):mI\preceq-\nabla^{2}\log p_{\theta}\preceq MI,~{}u_{\theta,j}\in\mathcal{U}_{\theta,j}\Big{\}},\\ &\mathcal{U}_{\theta,j}=\left\{u_{\theta,j}\in C^{2}_{0}(\mathbb{I}^{d_{N_{j}}}):\|u_{\theta,j}\|_{C^{2}}\leq R\right\}.\end{split}$ (4.5)
In practice, it can be implemented, for instance, by the Input Convex Neural Networks [1].
Theorem 4.1.
Given a $s$-sparse graph $G=(V,E)$ with ambient dimension $r$. Consider approximating the target distribution $p(x)\in\mathcal{P}_{1}(\mathbb{I}^{d})$ using localized score matching with the hypothesis class $\mathscr{P}$ defined above. Then with probability at least $1-\delta$, it holds that
$\max_{v\in V}W_{1}(p_{j},\hat{p}_{j})\leq CRN^{-1/4}\left(\log\delta^{-1}+\log b\right)^{1/4},$
where $\hat{p}:=p_{\hat{\theta}_{*}}$, $C$ is a constant depending only on $s,r,m,M$ and the local dimension $d_{\rm loc}$, thus independent of the total dimension $d$; and $b$ is the number of blocks, and is the only term of $O(d)$.
Proofs can be found in Appendix C. The theorem implies that sample complexity of the localized score mathching is bounded by $N_{\varepsilon}=O(\varepsilon^{-4}\log(d/\delta))$, with only logarithm dependence on the total dimension $d$, exactly characterizing that learning a localized distribution does not suffer from the curse of dimensionality. We also comment that this result does not achieve a minimax rate for estimating the target distribution $p$, as this is a corollary of the statistical estimation of the local score functions.
5 Application III: localized neural network surrogate
In this section, we consider the problem of constructing a surrogate for high dimensional distributions. It is ubiquitous that the target distribution is accessible, but is computationally expensive to evaluate. A common strategy is to construct a surrogate of the target distribution that is more computationally efficient, but still preserves the required accuracy. In this section, we investigate how the localized structure in the distribution helps to reduce the complexity of finding the surrogate.
As in the previous section, we assume the target distribution $p(x)$ has dependence graph $G=(V,E)$ that is $s$-sparse with ambient dimension $r$. We will construct a surrogate distribution $p_{\theta}(x)$ that has the same dependence graph $G$, and is easier to compute and sample from. There are various ways to construct the surrogate, e.g. vartional autoencoder [15], or the localized score matching in the previous section, but we do not specify the training method here.
In this section, we focus on the approximation capacity of the surrogate, and specifically consider using neural networks as the surrogate. We take the hypothesis class to be
$\mathscr{P}_{\rm NN}=\Big{\{}p_{\theta}(x)=\exp\left(\sum_{j\in V}u_{\theta,j}(x_{N_{j}})\right):mI\preceq-\nabla^{2}\log p_{\theta}\preceq MI,~{}u_{\theta,j}\in\Phi_{d_{N_{j}}}(L,W)\Big{\}},$ (5.1)
where we denote the fully connected neural network function class [28]
$\Phi_{d_{0}}(L,W)=\{\mathcal{N}_{\theta}:\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}|\theta=\{\mathcal{W}^{(l)},b^{(l)}\}_{l=1}^{L}\in\Theta_{d_{0}}(L,W)\}.$
Here $L,W$ are the depth and width of the neural function
$\mathcal{N}_{\theta}(x)=(\mathcal{W}^{(L)}\sigma(\cdot)+b^{(L)})\circ\cdots\circ(\mathcal{W}^{(1)}x+b^{(1)}),$
where $\sigma$ is the activation function. $\Theta_{d_{0}}(L,W)$ denotes the admissible parameter class
$\Theta_{d_{0}}(L,W)=$$\Big{\{}\{\mathcal{W}^{(l)},b^{(l)}\}_{l=1}^{L}:\mathcal{W}^{(1)}\in\mathbb{R}^{W\times d_{0}},~{}\mathcal{W}^{(l)}\in\mathbb{R}^{W\times W}~{}(1<l<L),$
$~{}\mathcal{W}^{(L)}\in\mathbb{R}^{1\times W};b^{(l)}\in\mathbb{R}^{W}~{}(1\leq l<L),~{}b^{(L)}\in\mathbb{R}\Big{\}}.$
In this section, we take the activation function to be the ReLU function $\sigma(x)=\max\{0,x\}$.
Note the complexity of the neural network is greatly reduced due to the sparse graphical structure. We show that
Theorem 5.1.
Given a $s$-sparse graph $G=(V,E)$ with ambient dimension $r$. Consider approximating the target distribution $p\in\mathscr{P}$ (4.5) using the neural network surrogate $p_{\theta}\in\mathscr{P}_{\rm NN}$ (5.1). Then there exists constants $C_{1},C_{2}$ depending only on the local dimension $d_{\rm loc}$, such that if we take the hyperparameters $L=C_{1}L_{0}\log L_{0}$ and $W=C_{2}W_{0}\log W_{0}$ for the NN class $\mathscr{P}_{\rm NN}$, then $\exists p_{\theta}\in\mathscr{P}_{\rm NN}$ s.t.
$\max_{j\in V}W_{1}(p_{j},p_{\theta,j})\leq CRL_{0}^{-2/d_{\rm loc}}W_{0}^{-2/d_{\rm loc}}.$
Here $C$ is a constant depending on $s,r,m,M$ and the local dimension $d_{\rm loc}$, thus independent of the total dimension $d$.
Proofs can be found in Appendix D. Similar to 4.1, the above theorem shows in the approximation perspective that learning the localized distribution does not suffer from the curse of dimensionality.
6 Proof on the gradient estimate of Stein equation
In this section, we illustrate the main steps to obtain gradient estimate of the Stein equation, which is the core of 2.2 and 2.7.
6.1 Proof of 2.2
Recall the Stein equation for the distribution $\pi$, index $i$ and $1$-Lipschitz function $\phi$ is
$\Delta u(x)+\nabla\log\pi(x)\cdot\nabla u(x)=\phi(x_{i}).$
The existence of its solution is standard in elliptic theory. We will use an explicit solution by Dynkin’s formula to derive the gradient estimate. Here the main technical tool is the stochastic analysis, which can cover all the three assumptions in this paper. This extends our previous method in [8] which mainly uses the PDE method.
Let $X_{t}^{x}=X_{t}^{x}(\omega)$ be the pathwise solution of the overdamped Langevin dynamics
$\mathrm{d}X_{t}^{x}(\omega)=\nabla\log\pi(X_{t}^{x}(\omega))\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t}(\omega),\quad X_{0}^{x}(\omega)=x.$ (6.1)
By Dynkin’s formula [18],
$\mathbb{E}u(X_{T}^{x})-u(x)=\mathbb{E}\int_{0}^{T}\left(\nabla\log\pi(X_{t}^{x})\cdot\nabla u(X_{t}^{x})+\Delta u(X_{t}^{x})\right)\mathrm{d}t=\int_{0}^{T}\mathbb{E}\phi(X_{t,i}^{x})\mathrm{d}t.$ (6.2)
Under any condition in 2.2, the distribution is strongly log-concave (see 2.3), and it is well-known that Law$(X_{t}^{x})$ converges to the equilibrium $\pi$ exponentially. This implies that the limit $T\rightarrow\infty$ exists for both sides in (6.2) (note $\int\phi(x_{i})\pi(x)\mathrm{d}x=0$), and the limit is
$\int u(x)\pi(x)\mathrm{d}x-u(x)=\int_{0}^{\infty}\mathbb{E}\phi(X_{t,i}^{x})\mathrm{d}t.$
Taking derivative w.r.t $x_{j}$ gives
$\nabla_{j}u(x)=-\int_{0}^{\infty}\mathbb{E}\left[\nabla_{j}X_{t,i}^{x}\cdot\nabla\phi(X_{t,i}^{x})\right]\mathrm{d}t.$
Here $\nabla_{j}X_{t,i}^{x}$ is the sample path partial derivative w.r.t $x_{j}$ and note $\omega$ is fixed. Note taking derivative on both sides is valid due to the exponential decay of $\nabla_{j}X_{t,i}^{x}$. Since $\phi$ is $1$-Lipschitz, we obtain
$\left|\nabla_{j}u(x)\right|\leq\int_{0}^{\infty}\mathbb{E}\left[\|\nabla_{j}X_{t,i}^{x}\|\left|\nabla\phi(X_{t,i}^{x})\right|\right]\mathrm{d}t\leq\mathbb{E}\int_{0}^{\infty}\|\nabla_{j}X_{t,i}^{x}\|\mathrm{d}t.$ (6.3)
So that it remains to control $\nabla_{j}X_{t,i}^{x}$. Taking derivative w.r.t. $x$ in (6.1) with $\omega$ fixed, we get
$\mathrm{d}\nabla X_{t}^{x}=-H_{t}\cdot\nabla X_{t}^{x}\mathrm{d}t,\quad H_{t}:=-\nabla^{2}\log\pi(X_{t}^{x}).$ (6.4)
where we denote for brevity $\nabla X_{t}^{x}:=\nabla_{x}X_{t}^{x}(\omega)\in\mathbb{R}^{d\times d}$. We then apply different assumptions in 2.2 to get the gradient estimate.
$\ell_{1}$-condition
Consider the partial derivative $\nabla_{j}X_{t}^{x}\in\mathbb{R}^{d\times d_{j}}$ in (6.4)
$\mathrm{d}\nabla_{j}X_{t}^{x}=-H_{t}\cdot\nabla_{j}X_{t}^{x}\mathrm{d}t.$
We aim to control the $2$-norm of the subblock $\nabla_{j}X_{t,k}^{x}\in\mathbb{R}^{d_{k}\times d_{j}}~{}(k=1,\dots b)$. Consider fixing a test vector $v_{j}\in\mathbb{R}^{d_{j}}$ s.t. $\left|v_{j}\right|=1$, and denote $g_{t}^{x}(k,j)=\nabla_{j}X_{t,k}^{x}\cdot v_{j}\in\mathbb{R}^{d_{k}}$, then
$\frac{\mathrm{d}}{\mathrm{d}t}g_{t}^{x}(k,j)=\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{j}X_{t,k}^{x}v_{j}=-H_{t}\cdot\nabla_{j}X_{t}^{x}v_{j}=-\sum_{l=1}^{b}H_{t}(k,l)\nabla_{j}X_{t,l}^{x}v_{j}=-\sum_{l=1}^{b}H_{t}(k,l)g_{t}^{x}(l,j).$
where we denote $H_{t}(k,l)\in\mathbb{R}^{d_{k}\times d_{l}}$ as the $(k,l)$-th subblock of $H_{t}$. Then by assumption,
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left|g_{t}^{x}(k,j)\right|^{2}=~{}$ $-\sum_{l=1}^{b}(g_{t}^{x}(k,j))^{\rm T}H_{t}(k,l)g_{t}^{x}(l,j)$
[…]
Notice $\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left|g_{t}^{x}(k,j)\right|^{2}=\left|g_{t}^{x}(k,j)\right|\frac{\mathrm{d}}{\mathrm{d}t}\left|g_{t}^{x}(k,j)\right|$, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\left|g_{t}^{x}(k,j)\right|\leq-M_{kk}\left|g_{t}^{x}(k,j)\right|+\sum_{l\neq k}M_{kl}\left|g_{t}^{x}(l,j)\right|.$
Such inequality holds for any index $j,k\in[b]$, test vector $v_{j}$ and initial condition $x$. For different indices $j\in[b]$, consider different initial conditions $x^{(j)}$, and denote the matrix $G_{t}\in\mathbb{R}^{b\times b}$ where
$G_{t}(k,j)=|g_{t}^{x^{(j)}}(k,j)|.$
Then the above inequality can be written compactly in a matrix form
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}\leq-\tilde{M}G_{t},\quad\tilde{M}_{ij}:=2M_{ii}\mathbf{1}_{i=j}-M_{ij}.$
Here $\leq$ is in the entrywise sense. Note the initial condition is $G_{0}=I_{b}$, since
$G_{0}(k,j)=|g_{0}^{x^{(j)}}(k,j)|=|\nabla_{x_{j}^{(j)}}X_{0,k}^{x^{(j)}}v_{j}|=|\nabla_{x_{j}^{(j)}}x_{k}^{(j)}v_{j}|=\delta_{jk}\left|v_{j}\right|=\delta_{jk}.$
By assumption, $\forall i,j\in[b],i\neq j$, $\tilde{M}_{ij}=-M_{ij}\leq 0$, and
$\sum_{j\neq i}|\tilde{M}_{ij}|+c=\sum_{j\neq i}M_{ij}+c\leq M_{ii}=\tilde{M}_{ii}.$
Thus we can apply A.1 and obtain
$\|G_{t}\|_{\infty}\leq\mathrm{e}^{-ct}\|G_{0}\|_{\infty}=\mathrm{e}^{-ct}~{}\Rightarrow~{}\max_{k}\sum_{j}|\nabla_{x_{j}^{(j)}}X_{t,k}^{x^{(j)}}v_{j}|\leq\mathrm{e}^{-ct}.$
Since $v_{j}$ is arbitrary, we obtain that
$\max_{k}\sum_{j}\|\nabla_{x_{j}^{(j)}}X_{t,k}^{x^{(j)}}\|\leq\mathrm{e}^{-ct}.$
Recall (6.3), this implies
$\sum_{j}|\nabla_{j}u(x^{(j)})|\leq~{}$ $\sum_{j}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\mathrm{d}t$
[…]
Now as $x^{(j)}$ is arbitrary, we obtain the gradient estimate
$\left\|\nabla u\right\|_{\infty,1}=\sum_{j}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq c^{-1}.$
$\ell_{2}$-condition
Consider $\nabla X_{t}^{x}$ in (6.4). Denote $G_{t}=\mathrm{e}^{mt}\nabla X_{t}^{x}$ and $\tilde{H}_{t}=H_{t}-mI$, then it holds that
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}=\mathrm{e}^{mt}\left(m\nabla X_{t}^{x}-H_{t}\nabla X_{t}^{x}\right)=-\tilde{H}_{t}G_{t},\quad G_{0}=\nabla X_{0}^{x}=I_{d}.$ (6.5)
By assumption, $0\preceq\tilde{H}_{t}\preceq(M-m)I$, and $\tilde{H}_{t}$ has bandwidth $B$. By A.2, it holds that
$\mathrm{e}^{mt}\|\nabla_{j}X_{t,i}^{x}\|=\|G_{t}(j,i)\|\leq\mathrm{e}^{-(M-m)t}\sum_{k=\lceil|i-j|/B\rceil}^{\infty}\frac{t^{k}(M-m)^{k}}{k!}.$
Note this estimate holds for any initial condition $x$. For different indices $j\in[b]$, consider different initial conditions $x^{(j)}$, and take summation over $j$, we obtain
$\sum_{j}~{}$ $\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\leq\mathrm{e}^{-mt}\mathrm{e}^{-(M-m)t}\sum_{j}\sum_{k=\lceil|i-j|/B\rceil}^{\infty}\frac{t^{k}(M-m)^{k}}{k!}$
[…]
Recall (6.3), this implies
$\sum_{j}|\nabla_{j}u(x^{(j)})|\leq~{}$ $\sum_{j}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\mathrm{d}t=\mathbb{E}\int_{0}^{\infty}\sum_{j}\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\mathrm{d}t$
[…]
Taking supremum over $x^{(j)}$, we obtain the gradient estimate
$\left\|\nabla u\right\|_{\infty,1}=\sum_{j}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq\frac{2BM}{m^{2}}.$
6.2 Proof of 2.7
The proof is similar to that of the $\ell_{2}$-condition case in the above section. For brevity, we only represent the different parts for the sparse graph treatment.
First note (6.3) is still valid, and we will use (6.5) to control $\nabla X_{t}^{x}$. For $G_{1}$ in (6.5), we need to modify the proof in A.2. Follow the notations therein, we can truncate the Dyson series as
$G_{1}=\exp(-M)\sum_{q=0}^{n}\frac{1}{q!}\hat{X}_{q}[X](1)+\exp(-M)\sum_{q=n+1}^{\infty}\frac{1}{q!}\hat{X}_{q}[X](1).$
Consider the off-diagonal entry $G_{1}(j,k)$. Take $n=d_{G}(j,k)-1$, then since all the path connecting $j$ and $k$ has length no less than $d_{G}(j,k)>n$, it must hold that
$\forall 1\leq q\leq n,\quad\left[(MI-H_{t_{1}})\cdots(MI-H_{t_{q}})\right](j,k)=0~{}\Rightarrow~{}\hat{X}_{q}[X](1)(j,k)=0,$
Therefore,
$G_{1}(j,k)=\exp(-M)\sum_{q=n+1}^{\infty}\frac{1}{q!}\hat{X}_{q}[X](1)(j,k).$
Note $\|\hat{X}_{q}[X](1)\|_{\rm op}\leq M^{q}$ by (A.2), we obtain
$\|G_{1}(j,k)\|\leq\exp(-M)\sum_{q=n+1}^{\infty}\frac{M^{q}}{q!}=\exp(-M)\sum_{q=d_{G}(j,k)}^{\infty}\frac{M^{q}}{q!}.$
By the same scaling argument as in the proof in A.2, we obtain
$\mathrm{e}^{mt}\|\nabla_{j}X_{t,k}^{x}\|=\|G_{t}(j,k)\|\leq\mathrm{e}^{-(M-m)t}\sum_{q=d_{G}(j,k)}^{\infty}\frac{t^{q}(M-m)^{q}}{q!}.$
The estimate holds for any initial condition $x$. For different vertices $j\in V$, consider different initial conditions $x^{(j)}$, and take summation over $j\in V$, we obtain
$\sum_{j\in V}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\leq~{}$ $\mathrm{e}^{-mt}\mathrm{e}^{-(M-m)t}\sum_{j\in V}\sum_{q=d_{G}(j,k)}^{\infty}\frac{t^{q}(M-m)^{q}}{q!}$
[…]
Here we use the sparsity condition $|N_{j}^{q}\backslash\{v\}|\leq sq^{r}$. From (6.3), we obtain
$\sum_{j\in V}|\nabla_{j}u(x^{(j)})|\leq~{}$ $\sum_{j\in V}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\mathrm{d}t=\mathbb{E}\int_{0}^{\infty}\sum_{j\in V}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\mathrm{d}t$
[…]
By A.4, it holds that (denote $\kappa=\frac{M}{m}$)
$\sum_{j\in V}|\nabla_{j}u(x^{(j)})|\leq~{}$ $\frac{1}{m}+\frac{s}{M}r!\left(\frac{m}{M}\right)^{-r-1}\left(1-\frac{m}{M}\right)$
[…]
Taking supremum over $x^{(j)}$, we obtain the gradient estimate.
6.3 Proof for 2.8
The proof is based on the proof of 2.4 and 2.2. And we only represent the different part for the multiple block case.
For any index set $I\subset[b]$, denote $d_{I}=\sum_{i\in I}d_{i}$. Then by definition,
$W_{1}(\pi_{I},\pi_{I}^{\prime})=\sup_{\phi_{0}\in{\rm Lip}_{1}(\mathbb{R}^{d_{I}})}\int\phi_{0}(x_{I})\left(\pi_{I}(x)-\pi_{I}^{\prime}(x)\right)\mathrm{d}x_{I}.$
Denote $\phi(x_{I})=\phi_{0}(x_{I})-\int\phi_{0}(x_{I})\pi^{\prime}(x_{I})\mathrm{d}x_{I}$, and let $u(x)$ solves the Stein equation
$\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)=\phi(x_{I}).$
Then using the same method as in 2.4, we obtain
$\int\phi_{0}(x_{I})\left(\pi_{I}(x)-\pi_{I}^{\prime}(x)\right)\mathrm{d}x_{I}=\sum_{j=1}^{b}\int\nabla_{j}u(x)\cdot\left(\nabla_{j}\log\pi^{\prime}(x)-\nabla_{j}\log\pi(x)\right)\pi(x)\mathrm{d}x.$
$\Rightarrow W_{1}(\pi_{I},\pi_{I}^{\prime})\leq\left\|\nabla u\right\|_{\infty,1}\cdot\max_{j}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.$
It remains to control $\left\|\nabla u\right\|_{\infty,1}$ for the new Stein equation. The method is still the same as in Section 6.1. Note here we need to replace (6.3) by
$\left|\nabla_{j}u(x)\right|\leq\int_{0}^{\infty}\mathbb{E}\left[\|\nabla_{j}X_{t,I}^{x}\|\left|\nabla\phi(X_{t,I}^{x})\right|\right]\mathrm{d}t\leq\mathbb{E}\int_{0}^{\infty}\|\nabla_{j}X_{t,I}^{x}\|\mathrm{d}t.$
Notice $\|\nabla_{j}X_{t,I}^{x}\|\leq\sum_{i\in I}\|\nabla_{j}X_{t,i}^{x}\|$, we obtain that
$\left\|\nabla u\right\|_{\infty,1}=\sum_{j\in[b]}\left|\nabla_{j}u(x)\right|\leq\sum_{j\in[b]}\mathbb{E}\int_{0}^{\infty}\sum_{i\in I}\|\nabla_{j}X_{t,i}^{x}\|\mathrm{d}t\leq\sum_{i\in I}\delta=\delta|I|.$
This completes the proof.
7 Conclusion
In this paper, we investigate the localized structure in high-dimensional distributions, and show how it can reduce the complexity of approximating and computing the localized distributions. We introduce the $\delta$-locality condition to quantify the degree of localization, and develop a $\ell_{\infty}$-approximation theory to get a finer control of the approximation error. This theory provides new insights into the localized structure in high-dimensional distributions, and establishes a theoretical foundation for localized methods. It suggests to exploit the localized structure in algorithm design to reduce the complexity of accessing these distributions. As examples, we apply the $\ell_{\infty}$-approximation theory to the localized LIS and localized score matching method, and show how the localized structure help to reduce the computational complexity. We also show the surrogate approximation of the localized distribution does not suffer from the curse of dimensionality.
Appendix A Lemmas
Lemma A.1.
Suppose $G_{t}\in\mathbb{R}_{+}^{b\times b}$ is a time-dependent nonnegative matrix satisfying
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}\leq-MG_{t},$
where $\leq$ is in the entrywise sense, and $M\in\mathbb{R}^{b\times b}$ is $\delta$-diagonal dominant with non-positive off-diagonal entries, i.e. $\forall i,j\in[b],i\neq j$,
$\sum_{j\neq i}|M_{ij}|+\delta\leq M_{ii};\quad M_{ij}\leq 0.$
Then for any $t\geq 0$, it holds
$\|G_{t}\|_{\infty}\leq\mathrm{e}^{-\delta t}\|G_{0}\|_{\infty}.$
Proof.
Denote $G_{t}^{\delta}:=\mathrm{e}^{\delta t}G_{t}$, then
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{\delta}=\mathrm{e}^{\delta t}\left(\frac{\mathrm{d}}{\mathrm{d}t}G_{t}+\delta G_{t}\right)\leq\mathrm{e}^{\delta t}\left(-MG_{t}+\delta G_{t}\right)=\left(-M+\delta I\right)G_{t}^{\delta}.$
Multiple both sides by $\mathbf{1}=(1,\dots,1)^{\rm T}\in\mathbb{R}^{b}$ from right,
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{\delta}\mathbf{1}\leq\left(-M+\delta I\right)G_{t}^{\delta}\mathbf{1}.$
This operator preserves the inequality since it is equivalent to taking summation over row indices. We claim that
$\forall t\geq 0,\quad|G_{t}^{\delta}\mathbf{1}|_{\infty}\leq|G^{\delta}_{0}\mathbf{1}|_{\infty},$ (A.1)
which is a reformulation of the conclusion $\left\|G_{t}\right\|_{\infty}\leq\left\|G_{0}\right\|_{\infty}$ by noticing $G_{t}^{\delta}\in\mathbb{R}_{+}^{b}$ and
$\mathrm{e}^{\delta t}\|G_{t}\|_{\infty}=\|G_{t}^{\delta}\|_{\infty}=\max_{i}\sum_{j}G_{t}^{\delta}(i,j)=|G_{t}^{\delta}\mathbf{1}|_{\infty}.$
We prove (A.1) by contraction. Suppose (A.1) is false, then $\exists s\geq 0$ and $i$ s.t.
$\frac{\mathrm{d}}{\mathrm{d}t}(G^{\delta}_{s}\mathbf{1})_{i}>0;\quad(G^{\delta}_{s}\mathbf{1})_{i}=|G^{\delta}_{s}\mathbf{1}|_{\infty}.$
On the other hand, notice by assumption on $M$,
$\frac{\mathrm{d}}{\mathrm{d}t}(G^{\delta}_{s}\mathbf{1})_{i}\leq~{}$ $\left(-M_{ii}+\delta\right)(G^{\delta}_{s}\mathbf{1})_{i}+\sum_{j\neq i}(-M_{ij})(G^{\delta}_{s}\mathbf{1})_{j}$
[…]
Contradiction. This proves our result. ∎
Lemma A.2.
Let $H_{t}\in\mathbb{R}^{d\times d}$ be a sequence of positive definite matrices with uniform block bandwidth $B>0$ and $2$-norm bound $M$, i.e.
$\forall|i-j|>B,\quad H_{t}(i,j)=0;\quad 0\preceq H_{t}\preceq MI_{d}.$
Consider the matrix ODE
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}=-H_{t}G_{t},\quad G_{0}=I_{d}.$
Then for any $t\geq 0$, it holds that
$\|G_{t}(i,j)\|\leq\exp(-tM)\sum_{k=\lceil|i-j|/B\rceil}^{\infty}\frac{(tM)^{k}}{k!}.$
Proof.
I. Scaling. First notice it suffices to consider the case $t=1$. For the general case $t=t_{0}$, consider the following scaling argument. Let $G_{t}$ be the solution, then $G_{t_{0}s}$ solves
$\frac{\mathrm{d}}{\mathrm{d}s}G_{t_{0}s}=-t_{0}H_{t_{0}s}G_{t_{0}s},\quad G_{0}=I_{d}.$
If the theorem holds for $t=1$, we obtain that at $s=1$ (notice $\left\|t_{0}H_{t_{0}s}\right\|\leq Mt_{0}$),
$\|G_{t_{0}}(i,j)\|\leq\exp(-Mt_{0})\sum_{k=\lceil|i-j|/B\rceil}^{\infty}\frac{(Mt_{0})^{k}}{k!}.$
Next we consider the case $t=1$.
II. Dyson series [7]. By variation of constants, we have
$G_{t}=I_{d}-\int_{0}^{t}H_{s}G_{s}\mathrm{d}s.$
Applying this identity recursively, we obtain
$G_{t}=~{}$ $I_{d}-\int_{0}^{t}H_{s}\left(I_{d}-\int_{0}^{s}H_{u}G_{u}\mathrm{d}u\right)\mathrm{d}s$
[…]
$+(-1)^{N}\int_{[0,t]^{N}}H_{t_{N}}\cdots H_{t_{1}}G_{t_{1}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{N}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{N}.$
For simplicity, denote
$X_{0}(t)=I_{d},\quad X_{n}(t):=\frac{n!}{t^{n}}\int_{[0,t]^{n}}H_{t_{n}}\cdots H_{t_{1}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n},\quad n\geq 1.$
$R_{N}(t)=\int_{[0,t]^{N}}H_{t_{N}}\cdots H_{t_{1}}G_{t_{1}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{N}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{N}.$
Notice $X_{n}(t)$ is the average of “$H^{n}$” in $[0,t]$. Then
$G_{t}=\sum_{n=0}^{N-1}(-1)^{n}\frac{t^{n}}{n!}X_{n}(t)+R_{N}(t).$
Now we prove that $\lim_{N\rightarrow\infty}R_{N}(t)=0$. First notice $\left\|G_{t}\right\|_{2}\leq 1$, since
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{\rm T}G_{t}=-2G_{t}^{\rm T}H_{t}G_{t}~{}\Rightarrow~{}G_{t}^{\rm T}G_{t}=I_{d}-\int_{0}^{t}G_{s}^{\rm T}H_{s}G_{s}\mathrm{d}s\preceq I_{d}.$
So that as $N\rightarrow\infty$,
$\left\|R_{N}(t)\right\|_{2}\leq~{}$ $\int_{[0,t]^{N}}\left\|H_{t_{N}}\right\|_{2}\cdots\left\|H_{t_{1}}\right\|_{2}\left\|G_{t_{1}}\right\|_{2}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{N}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}$
[…]
This proves that the Dyson series converges, and
$G_{t}=\sum_{n=0}^{\infty}(-1)^{n}\frac{t^{n}}{n!}X_{n}(t).$
III. Representation of polynomials. Denote the matrix process space
$\mathscr{X}=\text{span}\{X_{n}(t),n\geq 0\}=\overline{\Big{\{}\sum_{k=0}^{n}a_{k}X_{n}(t):a_{k}\in\mathbb{C}\Big{\}}}.$
We define the representation of any polynomial $P$ in $\mathscr{X}$ as
$P[X](t)=\sum_{k=0}^{n}a_{k}X_{k}(t),\quad\text{if }P(x)=\sum_{k=0}^{n}a_{k}x^{k}.$
In A.3 we show that the representation has an equivalent definition:
$P\{X\}(t)=\frac{a_{n}}{t^{n}}\sum_{\sigma\in S_{n}}\int_{0}^{t}(H_{t_{1}}-x_{\sigma_{1}}I)\int_{0}^{t_{1}}(H_{t_{2}}-x_{\sigma_{2}}I)\cdots\int_{0}^{t_{n-1}}(H_{t_{n}}-x_{\sigma_{n}}I)\mathrm{d}t_{n}\cdots\mathrm{d}t_{2}\mathrm{d}t_{1}.$
if $\text{if }P(x)=a_{n}\prod_{k=1}^{n}(x-x_{k}).$
Here $S_{n}$ is the set of all permutations of $[n]$. We use $P\{X\}$ to distinguish different ways of definitions. But they are in fact equivalent.
IV. Banded matrix approximation. By Taylor expansion,
$\exp(-x)=\exp(-M)\sum_{n=0}^{\infty}\frac{(M-x)^{n}}{n!}.$
Represent the series in $\mathscr{X}$, we obtain (denote $\hat{X}_{n}(x)=(M-x)^{n}$)
$G_{1}=\exp(-M)\sum_{n=0}^{\infty}\frac{1}{n!}\hat{X}_{n}[X](1).$
Consider its finite truncation
$G_{1}=\exp(-M)\sum_{k=0}^{n}\frac{1}{k!}\hat{X}_{k}[X](1)+\exp(-M)\sum_{k=n+1}^{\infty}\frac{1}{k!}\hat{X}_{k}[X](1).$
Notice by the alternative representation,
$\hat{X}_{n}[X](1)=\sum_{\sigma\in S_{n}}\int_{[0,1]^{n}}(MI-H_{t_{1}})\cdots(MI-H_{t_{n}})\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}.$
Since $0\preceq H_{t_{i}}\preceq MI$, it holds that $\left\|MI-H_{t_{i}}\right\|_{\rm op}\leq M$, and thus
$\|\hat{X}_{n}[X](1)\|_{\rm op}\leq M^{n}n!\int_{[0,1]^{n}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}=M^{n}.$ (A.2)
Since $H_{t}$ has block bandwidth $B$, the block bandwidth of $\hat{X}_{k}[X]$ is at most $kB$. So that if $|i-j|>nB\geq 0$, $\hat{X}_{k}[X](i,j)=0$ if $k\leq n$, and thus
$\|G_{1}(i,j)\|=~{}$ $\left\|\exp(-M)\sum_{k=n+1}^{\infty}\frac{1}{k!}\hat{X}_{k}[X](1)(i,j)\right\|$
[…]
This verifies the case $i\neq j$. For $i=j$, the result follows from $\|G_{1}(i,i)\|\leq\left\|G_{1}\right\|\leq 1$. ∎
Lemma A.3.
The two representations of polynomials in $\mathscr{X}$ are equivalent, i.e.
$P[X](t)=P\{X\}(t),\quad\forall P\in\mathbb{P}.$
Proof.
We prove by induction. First note $n=1$ is obvious,
$(a_{1}(x-x_{1}))\{X\}(t)=\frac{a_{1}}{t}\int_{0}^{t}(H_{t_{1}}-x_{1}I)\mathrm{d}t_{1}=a_{1}(X_{1}(t)-x_{1}I_{d})=(a_{1}(x-x_{1}))[X](t).$
Now consider $n\geq 2$. Notice $P[X](t),P\{X\}(t)$ can both be viewed as mulitlinear maps on $x_{1},\dots,x_{n}$. Compute the partial derivative w.r.t. $x_{n}$,
$\nabla_{x_{n}}\left(P\{X\}(t)\right)$
[…]
Here the first equality follows by discussion on the position of $x_{n}$. The third equality follows by integrating the variable $t_{k}$ and notice the constraint $t_{k+1}\leq t_{k}\leq t_{k-1}$. The forth equality follows by relabeling the index. The last equality follows by induction hypothesis. By symmetry, the relation holds for any $x_{i}$:
$\nabla_{x_{i}}\left(P\{X\}(t)\right)=-\left(a_{n}\prod_{j\neq i}(x-x_{j})\right)[X](t)=(\nabla_{x_{i}}P)[X](t),$
since $\nabla_{x_{i}}P(x)=-a_{n}\prod_{j\neq i}(x-x_{j})$. Now notice the representation $P[X]$ is linear in the coefficients of $P$, it is direct to verify
$\nabla_{x_{i}}(P[X](t))=(\nabla_{x_{i}}P)[X](t)=\nabla_{x_{i}}\left(P\{X\}(t)\right).$
Finally notice when $x_{1}=\cdots=x_{n}=0$,
$x^{n}\{X\}(t)=\frac{n!}{t^{n}}\int_{0}^{t}H_{t_{1}}\int_{0}^{t_{1}}H_{t_{2}}\cdots\int_{0}^{t_{n-1}}H_{t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}=X_{n}(t)=x^{n}\{X\}(t).$
Now the two multi-linear maps agree on one point, and also all the partial derivatives. So that they must be identical. This shows the equivalence holds for $n$, and by induction, it holds for all $n$. ∎
Lemma A.4.
For any $t\geq 0$ and $x\in(0,1)$, it holds that
$\sum_{k\geq 1}k^{t}(1-x)^{k}<2\Gamma(t+1)x^{-t-1}(1-x).$
When $t\in\mathbb{N}$, the factor $2$ can be omitted, i.e. $\sum_{k\geq 1}k^{t}(1-x)^{k}\leq t!x^{-t-1}(1-x)$.
Proof.
We prove by induction. For $t=0$, it holds that
$\sum_{k\geq 1}(1-x)^{k}=\frac{1-x}{x}.$
For $t\in(0,1)$, first notice by Abel transformation,
$\sum_{k\geq 1}k^{t}(1-x)^{k}=x^{-1}\sum_{k\geq 1}k^{t}\left[(1-x)^{k}-(1-x)^{k+1}\right]=x^{-1}\sum_{k\geq 1}\left(k^{t}-(k-1)^{t}\right)(1-x)^{k}.$
Therefore, since $k^{t}-(k-1)^{t}\leq t(k-1)^{t-1}$ when $t\in(0,1)$ and $k\geq 2$,
$\sum_{k\geq 1}k^{t}(1-x)^{k}\leq~{}$ $x^{-1}\left[(1-x)+\sum_{k\geq 2}t(k-1)^{t-1}(1-x)^{k}\right]$
[…]
Here we use $t\Gamma(t)=\Gamma(t+1)$ and $\Gamma(t+1)x^{-t}>1$ in the last step. This verifies the case $t\in[0,1)$. Suppose the inequality holds for $t-1\geq 0$, then using the same methods,
$\sum_{k\geq 1}k^{t}(1-x)^{k}=~{}$ $x^{-1}\sum_{k\geq 1}\left(k^{t}-(k-1)^{t}\right)(1-x)^{k}\leq x^{-1}\sum_{k\geq 1}tk^{t-1}(1-x)^{k}$
[…]
Here the first inequality follows by the elementary inequality $k^{t}-(k-1)^{t}\leq tk^{t-1}$ when $t\geq 1$, and the second inequality follows by induction hypothesis. The finer inequality for $t\in\mathbb{N}$ can be obtained similarly. This completes the proof. ∎
Appendix B Proofs for 3.2
By 3.1 and 2.7, denote $\delta=\frac{sr!\kappa^{r}}{m}$,
$\max_{j\in V}W_{1}(\pi_{j},\pi_{r,j})\leq\delta\cdot\max_{j\in V}\left\|\nabla_{j}\log\pi-\nabla_{j}\log\pi_{r}\right\|_{L^{1}(\pi_{r})}.$
Denote $x:=C^{-1/2}\tilde{x}$. By the chain rule,
$\nabla_{j}=\sum_{k\in V}\frac{\partial\tilde{x}_{k}}{\partial x_{j}}\cdot\tilde{\nabla}_{k}=\sum_{k\in V}C^{1/2}(j,k)\tilde{\nabla}_{k}.$
So that by B.1 (see below),
$\max_{j}\int|\nabla_{j}\log\pi(x)-\nabla_{j}\log\pi_{r}(x)|\pi_{r}(x)\mathrm{d}x$
[…]
So that one obtain (denote $C_{\pi}=m^{-3/2}(sr!\kappa^{r})^{2}$)
$\max_{j}W_{1}(\pi_{j},\pi_{r,j})\leq C_{\pi}\max_{k}\int|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})|\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}.$ (B.1)
Next we control the right hand side of (B.1). Note $\tilde{\pi}\propto\tilde{l}\tilde{\pi}_{0}$ and $\tilde{\pi}_{r}\propto\tilde{l}_{r}\tilde{\pi}_{0}$, it holds that
~$\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})=\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{l}_{r}(\tilde{x}_{r}),$
Due to the orthogonality of $\tilde{x}_{r}$ and $\tilde{x}_{\bot}$, one can decompose
|~$|\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2}=|\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})|^{2}+|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2},$
where we denote ~$\tilde{\nabla}_{r,k}=\nabla_{\tilde{x}_{r,k}},\tilde{\nabla}_{\bot,k}=\nabla_{\tilde{x}_{\bot,k}}$ and use ~$\tilde{\nabla}_{\bot,k}\log\tilde{l}_{r}(\tilde{x}_{r})=0$ as $\tilde{l}_{r}(\tilde{x}_{r})$ does not depend on $\tilde{x}_{\bot}$. Therefore,
|~$\begin{split}\int|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}=\int|\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}&\\ +\int|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}=:~{}&\text{I}+\text{II}.\end{split}$ (B.2)
The first term is straightforward to control. Notice
$\begin{split}\text{I}=~{}&\int\text{\rm tr}\left[\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x}))^{\rm T}\right]\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\\ =~{}&\text{\rm tr}\left[\int\tilde{P}_{\bot,k}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x}))^{\rm T}\tilde{P}_{\bot,k}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\right]\\ =~{}&\text{\rm tr}\left[\tilde{P}_{\bot,k}\widetilde{\mathcal{G}}_{k}\tilde{P}_{\bot,k}\right]=\mathcal{R}(\widetilde{\mathcal{G}}_{k},\widetilde{\mathcal{X}}_{r,k}).\end{split}$ (B.3)
For the second term, first notice $\tilde{\pi}_{r}$ can be decomposed as
$\tilde{\pi}_{r}(\tilde{x})=\tilde{Z}_{r}^{-1}\tilde{l}_{r}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x})=\tilde{Z}_{r}^{-1}\tilde{l}_{r}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x}_{\bot})=\tilde{\pi}_{r}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x}_{\bot}),$
where $\tilde{\pi}_{r}(\tilde{x}_{r})=\tilde{Z}_{r}^{-1}\tilde{l}_{r}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x}_{r})$ denotes the marginal distribution of $\tilde{\pi}_{r}$ on $\tilde{x}_{r}$. So that
$\text{II}=\int\left(\int|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}\right)\tilde{\pi}_{r}(\tilde{x}_{r})\mathrm{d}\tilde{x}_{r}.$
Take ~$\tilde{\nabla}_{r,k}$ in (3.4), we have
~$\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})=\int\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}.$
So that ~$\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})$ is the average of ~$\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})$ with respect to $\tilde{\pi}_{0}(\tilde{x}_{\bot})$. Therefore, we can apply the Poincaré inequality for the function ~$\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})$ with reference measure $\tilde{\pi}_{0}(\tilde{x}_{\bot})$ (which is a standard Gaussian) to obtain
|~$\int|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})|^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}$
[…]
Note one can relax
~$\|\tilde{\nabla}_{\bot,j}\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})\|_{\rm F}^{2}=~{}$ $\text{\rm tr}\left[\tilde{\nabla}_{\bot,j}\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{r,k}\tilde{\nabla}_{\bot,j}\log\tilde{l}(\tilde{x})\right]$
[…]
Therefore,
$\begin{split}\text{II}\leq~{}&\int\left(\int\sum_{j\in V}\|\tilde{\nabla}_{\bot,j}\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})\|_{\rm F}^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}\right)\tilde{\pi}_{r}(\tilde{x}_{r})\mathrm{d}\tilde{x}_{r}\\ \leq~{}&\int\sum_{j\in V}\text{\rm tr}\left[\tilde{P}_{\bot,j}\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{k}\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})\tilde{P}_{\bot,j}\right]\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\\ =~{}&\sum_{j\in V}\text{\rm tr}\left[\tilde{P}_{\bot,j}\widetilde{\mathcal{H}}_{j,k}\tilde{P}_{\bot,j}\right]=\sum_{v\in V}\mathcal{R}(\widetilde{\mathcal{H}}_{j,k},\widetilde{\mathcal{X}}_{r,j}).\end{split}$ (B.4)
Combine (B.1)-(B.4), we obtain
$\max_{j}W_{1}(\pi_{j},\pi_{r,j})\leq~{}$ $C_{\pi}\max_{k}\left(\int|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\right)^{1/2}$
[…]
where we use Cauchy’s inequality in the first step.
B.1 Lemma
Lemma B.1.
Under 3.1, it holds that (denote $\kappa=\frac{M}{m}$)
$\max_{j\in V}\sum_{k\in V}\|C^{1/2}(j,k)\|\leq m^{-1/2}sr!\kappa^{r}.$ (B.5)
Proof.
By Taylor expansion,
$x^{-1/2}=M^{-1/2}\left(1+\sum_{q=1}^{\infty}\frac{(2q-1)!!}{(2q)!!}\left(1-\frac{x}{M}\right)^{q}\right),\quad x\in[m,M].$
For $j\neq k$, let $n=d_{G}(j,k)-1$, plug in $x=C^{-1}$ and notice $mI\preceq C^{-1}\preceq MI$, we obtain
$C^{1/2}=P_{n}(C^{-1})+M^{-1/2}\sum_{q=n+1}^{\infty}\frac{(2q-1)!!}{(2q)!!}\left(I-\frac{C^{-1}}{M}\right)^{q}.$
where $P_{n}\in\mathbb{P}_{n}$. By assumption, $C^{-1}(u,v)=0$ if $v\notin N_{u}$. Since $d_{G}(j,k)=n+1$, it holds
$P_{n}(C^{-1})(j,k)=0,$
as all the path connecting $j$ and $k$ has length no less than $d_{G}(j,k)>n$. Therefore,
$\|C^{1/2}(j,k)\|=~{}$ $M^{-1/2}\left\|\sum_{q=n+1}^{\infty}\frac{(2q-1)!!}{(2q)!!}\left(I-\frac{C^{-1}}{M}\right)^{k}(j,k)\right\|$
[…]
For the coefficients, we bound
$\frac{(2q-1)!!}{(2q)!!}=~{}$ $\frac{1}{2}\prod_{j=2}^{q}\left(1-\frac{1}{2j}\right)\leq\frac{1}{2}\prod_{j=2}^{q}\exp\left(-\frac{1}{2j}\right)=\frac{1}{2}\exp\left(-\frac{1}{2}\sum_{j=2}^{q}\frac{1}{j}\right)$
[…]
Since $mI\preceq C^{-1}\preceq MI$, it holds that
$0\preceq\left(I-\frac{C^{-1}}{M}\right)^{q}\preceq\left(1-\frac{m}{M}\right)^{q}I~{}\Rightarrow~{}\left\|\left(I-\frac{C^{-1}}{M}\right)^{q}\right\|_{\rm op}\leq(1-\kappa^{-1})^{q}.$
For diagonal entry, $\|C^{1/2}(j,j)\|\leq\left\|C^{1/2}\right\|\leq m^{-1/2}$, so that
$\sum_{k\in V}\|C^{1/2}(j,k)\|=~{}$ $\|C^{1/2}(j,j)\|+\sum_{k\neq j}\|C^{1/2}(j,k)\|$
[…]
Here the second inequality follows the sparsity assumption, the third inequality follows A.4, and the last inequality follows from $\Gamma\left(r+\frac{1}{2}\right)\leq r!$ and $sr!\kappa^{r-1}\geq 1$. This completes the proof. ∎
Appendix C Proof of 4.1
By 2.7, $p_{\theta}\in\mathscr{P}$ is $\delta$-localized with $\delta=\frac{sr!\kappa^{r}}{m}$, and
$\begin{split}\max_{v\in V}W_{1}(p_{j},\hat{p}_{j})\leq~{}&\delta\cdot\max_{j\in V}\left\|\nabla_{j}\log p-\nabla_{j}\log\hat{p}\right\|_{L^{1}(p)}\\ \leq~{}&\delta\cdot\left(\max_{j\in V}\mathbb{E}_{X\sim p}|s(X)-s_{\hat{\theta}_{*},v}(X)|^{2}\right)^{1/2}=\delta\cdot J_{\rm loc}^{1/2}(\hat{\theta}_{*}).\end{split}$ (C.1)
Using optimality of $\hat{\theta}_{*}$ and $J_{\rm loc}(\theta_{*})=0$,
$J_{\rm loc}(\hat{\theta}_{*})=~{}$ $\left(J_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\hat{\theta}_{*})\right)+\left(\hat{J}_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\theta_{*})\right)+\left(\hat{J}_{\rm loc}(\theta_{*})-J_{\rm loc}(\theta_{*})\right)$
[…]
Given data $X_{1},\dots,X_{N}$, denote for simplicity
$\mathbb{P}_{N}f=\frac{1}{N}\sum_{i=1}^{N}f(X_{i}),\quad f_{\theta,j}(x)=2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(x)\right)+\left|s_{\theta,j}(x)\right|^{2}.$
Then $\hat{J}_{j}(\theta)-J_{j}(\theta)=(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta,j}$. For any $\theta,\phi\in\Theta$, it holds that
$|f_{\theta,j}-f_{\phi,j}$ $|\leq 2\left|\Delta_{jj}\log p_{\theta}-\Delta_{jj}\log p_{\phi}\right|+\left|\left|\nabla_{j}\log p_{\theta}\right|^{2}-\left|\nabla_{j}\log p_{\phi}\right|^{2}\right|$
[…]
where we use $|N_{j}|\leq s+1$ and $\left|\nabla_{j}u_{\theta,k}(x)\right|\leq R$.
Consider the function space
$\mathcal{U}_{\theta}:=\{u_{\theta}=(u_{\theta,j})_{j\in V}:u_{\theta,j}\in\mathcal{U}_{\theta,j}\}.$
Equip $\mathcal{U}_{\theta}$ with norm
$\left\|u_{\theta}\right\|_{\mathcal{U}}:=\max_{j\in V}\left\|u_{\theta,j}\right\|_{C^{2}}.$
By definition, $\sup_{u_{\theta}\in\mathcal{U}_{\theta}}\left\|u_{\theta}\right\|_{\mathcal{U}}\leq R$. Consider a $\varepsilon$-net $\mathcal{U}_{\varepsilon}=\{u_{i}\}_{n=1}^{N_{\varepsilon}}\subset\mathcal{U}_{\theta}$ of $\mathcal{U}_{\theta}$ with respect to $\left\|\cdot\right\|_{\mathcal{U}}$, i.e. $\forall u_{\theta}\in\mathcal{U}_{\theta}$, there exists some $u_{i}\in\mathcal{U}_{\varepsilon}$ s.t. $\left\|u_{\theta}-u_{i}\right\|_{\mathcal{U}}\leq\varepsilon$. Then
$\left\|f_{\theta,j}-f_{\theta_{i},j}\right\|_{L^{\infty}}\leq 2(s+1)R\sum_{k\in N_{j}}\left\|u_{\theta,k}-u_{\theta_{i},k}\right\|_{C^{2}}\leq 2(s+1)^{2}R\varepsilon.$
Here we take $\theta_{i}$ to be the parameter corresponding to $u_{i}$. Therefore,
$|\hat{J}_{j}(\theta)-J_{j}(\theta)|=~{}$ $|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta,j}|\leq|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|+|(\mathbb{P}_{N}-\mathbb{E}_{p})\left(f_{\theta,j}-f_{\theta_{i},j}\right)|$
[…]
This further implies that
$\max_{\theta\in\Theta}\max_{j\in V}|\hat{J}_{j}(\theta)-J_{j}(\theta)|\leq\max_{1\leq i\leq N_{\varepsilon}}\max_{j\in V}|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|+2(s+1)^{2}R\varepsilon.$
Next we give a bound on $N_{\varepsilon}$. It is a classical result (see for instance [24]) that the covering number of the unit ball $B_{1}(C^{2}(\mathbb{I}^{d_{N_{j}}})):=\{u\in C^{2}(\mathbb{I}^{d_{N_{j}}}):\left\|u\right\|_{C^{2}}\leq 1\}$ is bounded by
$\log N(\varepsilon,B_{1}(C^{2}(\mathbb{I}^{d_{N_{j}}})),L^{\infty})\leq C_{d_{N_{j}}}\varepsilon^{-d_{N_{j}}/2}.$
Here $\mathcal{U}_{\theta}$ is the product space of $C^{2}(\mathbb{I}^{d_{N_{j}}})$ equipped with the $\ell_{\infty}$-norm, so that
$\log N_{\varepsilon}\leq\sum_{j\in V}\log N(\varepsilon,B_{1}(C^{2}(\mathbb{I}^{d_{N_{j}}})),L^{\infty})\leq bC_{d_{\rm loc}}\varepsilon^{-d_{\rm loc}/2}.$
Recall here $d_{\rm loc}=\max_{j}d_{N_{j}}$. Notice $f_{\theta,j}$ is uniformly bounded by
$|f_{\theta,j}|\leq 2(s+1)R\sum_{k\in N_{j}}\left\|u_{\theta,k}\right\|_{C^{2}}\leq 2(s+1)^{2}R.$
By Hoeffding’s inequality,
$\mathbb{P}\left[|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq\mathrm{e}^{-Nt^{2}/2(s+1)^{4}R^{2}},\quad\forall t>0.$
So that by maximal inequality,
$\mathbb{P}\left[\max_{1\leq i\leq N_{\varepsilon}}\max_{j\in V}|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq N_{\varepsilon}b\cdot\mathrm{e}^{-Nt^{2}/2(s+1)^{4}R^{2}}.$
Take $t=(s+1)^{2}R\varepsilon$, then
$\mathbb{P}\left[\max_{\theta\in\Theta}\max_{j\in V}|\hat{J}_{j}(\theta)-J_{j}(\theta)|\geq 3t\right]\leq\mathbb{P}\left[\max_{1\leq i\leq N_{\varepsilon}}\max_{j\in V}|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq N_{\varepsilon}b\cdot\mathrm{e}^{-N\varepsilon^{2}/2}.$
Combine the above estimate, we obtain
$\mathbb{P}\left[J_{\rm loc}(\hat{\theta}_{*})\geq 6(s+1)^{2}R^{2}\varepsilon\right]\leq b^{2}C_{d_{\rm loc}}\varepsilon^{-d_{\rm loc}/2}\mathrm{e}^{-N\varepsilon^{2}/2}.$
In order to make the right hand side less than some $\delta>0$, one can take
$\varepsilon^{2}=\frac{C}{N}\left(\log\delta^{-1}+\log b\right),$
where $C$ is some constant depending only on $d_{\rm loc}$. So that with probability at least $1-\delta$,
$J_{\rm loc}(\hat{\theta}_{*})\leq C^{\prime}N^{-1/2}\sqrt{\log\delta^{-1}+\log b}.$
The conclusion follows by plugging in the bound (C.1) on $\max_{j}W_{1}(p_{j},\hat{p}_{j})$.
Appendix D Proof for 5.1
The proof is based on the following approximation result adapted from Theorem 3 in [28]:
Proposition D.1.
For any $u\in W^{2,\infty}(\mathbb{I}^{d_{0}})$ and $\left\|u\right\|_{W^{2,\infty}}\leq 1$, there exists a neural network $u_{\theta}\in\Phi_{d_{0}}(L,W)$ where $L=C_{1}L_{0}\log L_{0}$ and $W=C_{2}W_{0}\log W_{0}$ such that
$\left\|u-u_{\theta}\right\|_{W^{1,\infty}(\mathbb{I}^{d_{0}})}\leq C_{3}L_{0}^{-2/d_{0}}W_{0}^{-2/d_{0}}.$
Here $C_{1},C_{2},C_{3}$ are constants depending only on $d_{0}$.
By 2.7, for $p_{\theta}\in\mathscr{P}_{\rm NN}$, it holds that (denote $\delta=\frac{sr!\kappa^{r}}{m}$)
$\max_{j\in V}W_{1}(p_{j},p_{\theta,j})\leq\delta\cdot\max_{k\in V}\left\|\nabla_{k}\log p-\nabla_{k}\log p_{\theta}\right\|_{L^{1}(p)}.$
By the sprase graphical structure,
$\nabla_{k}\log p-\nabla_{k}\log p_{\theta}=\sum_{v\in N_{k}}\left(\nabla_{k}u_{j}(x_{N_{j}})-\nabla_{k}u_{\theta,j}(x_{N_{j}})\right).$
Therefore, since $G$ is $s$-sparse,
$\left\|\nabla_{k}\log p-\nabla_{k}\log p_{\theta}\right\|_{L^{1}(p)}\leq$ $\sum_{v\in N_{k}}\left\|\nabla_{k}u_{j}-\nabla_{k}u_{\theta,j}\right\|_{L^{1}(p)}$
$\leq(s+1)\max_{j\in V}$ $\left\|\nabla_{k}u_{j}-\nabla_{k}u_{\theta,j}\right\|_{L^{\infty}}\leq(s+1)\max_{j\in V}\left\|u_{j}-u_{\theta,j}\right\|_{W^{1,\infty}}.$
We obtain
$\max_{j\in V}W_{1}(p_{j},p_{\theta,j})\leq\delta(s+1)\cdot\max_{j\in V}\left\|u_{j}-u_{\theta,j}\right\|_{W^{1,\infty}}.$
By D.1, we can take $L=C_{1}L_{0}\log L_{0}$ and $W=C_{2}W_{0}\log W_{0}$ in $\mathscr{P}_{\rm NN}$, so that for any $j\in V$, there exists $u_{\theta,j}\in\Phi_{d_{N_{j}}}(L,W)$ s.t. (notice we take $\left\|u_{j}\right\|_{W^{2,\infty}}\leq R$)
$\left\|u_{j}-u_{\theta,j}\right\|_{W^{1,\infty}}\leq C_{3}RL_{0}^{-2/d_{N_{j}}}W_{0}^{-2/d_{N_{j}}}\leq C_{3}RL_{0}^{-2/d_{\rm loc}}W_{0}^{-2/d_{\rm loc}}.$
Here $d_{\rm loc}=\max_{j\in V}d_{N_{j}}$, and the constant $C_{1},C_{2},C_{3}$ depend only on $d_{\rm loc}$. This completes the proof.