### v2

# Stein’s method for marginals on large graphical models

Abstract
Many spatial models exhibit locality structures that effectively reduce their intrinsic dimensionality, enabling efficient approximation and sampling of high-dimensional distributions. However, existing approximation techniques mainly focus on joint distributions, and do not guarantee accuracy for low-dimensional marginals. By leveraging the locality structures, we establish a dimension independent uniform error bound for the marginals of approximate distributions. Inspired by the Stein’s method, we introduce a novel $\delta$-locality condition that quantifies the locality in distributions, and link it to the structural assumptions such as the sparse graphical models. The theoretical guarantee motivates the localization of existing sampling methods, as we illustrate through the localized likelihood-informed subspace method and localized score matching. We show that by leveraging the locality structure, these methods greatly reduce the sample complexity and computational cost via localized and parallel implementations.
1 Introduction
Approximation and sampling of high-dimensional distributions are fundamental challenges in computational mathematics, statistics and machine learning [27, 18, 23]. A general strategy to mitigate these challenges is to exploit low-dimensional structures in real-world distributions, such as the sparsity [22] and the manifold assumption [16]. In this paper, we study a type of locality structure [4, 35] that models the sparse dependence between components in high-dimensional distributions. This structure arises naturally in spatial models, and has been successfully applied in various fields such as spatial statistics [4, 44, 13], data assimilation [25, 21, 35], and quantum mechanics [28, 3, 6]. Recently, there has been a surging interest for its applications in distribution sampling problems [46, 41, 17, 19, 8]. The goal of this paper is to introduce a quantification of locality structure in high-dimensional distributions, and to develop efficient approximation and sampling methods that exploit this structure to achieve scalability in high dimensions.
Mathematically speaking, the accuracy of various approximation and sampling algorithms are guaranteed by distribution inequalities [34, 20]. For example, denote $1$-Wasserstein distance as $\mathsf{W}_{1}$, use Stein’s Lemma, and one can estabilish an Otto-Villani type inequality [34]
$\mathsf{W}_{1}(\pi,\pi^{\prime})\leq C_{\pi}\mathbb{E}_{\pi^{\prime}}\left[\|\nabla\log\pi-\nabla\log\pi^{\prime}\|\right],$ (1.1)
where $C_{\pi}$ is some constant relates to the log-concavity of $\pi$. Note the right hand side of (1.1) is the difference of the gradient of the log density $\nabla\log\pi$ (the score function), which can be used as a loss function in many distribution estimation and approximation methods, including the score matching and diffusion models [26, 37, 23].
However, (1.1) can be awarkward to use when the distributions are high dimensional. Since the score functions are $d$ dimensional vectors, the right hand side of (1.1) usually grows like $\sqrt{d}$. This may suggest it is necessary to use more training data or finer computational schemes for larger models. The main objective of this paper is to break such pessimistic view by leveraging locality structure. The highlight of the analysis is that the approximation error of low-dimensional marginals can be independent of the ambient dimension, which facilitates efficient approximation and sampling methods [32, 41, 17, 8]. We introduce a $\delta$-locality condition (2.1) to quantify the locality structure, and establish a dimension independent bound for the marginal error of approximate distributions (3.1):
$\max_{i\in[d]}\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})\leq\delta\cdot\max_{j\in[d]}\int\left\|\nabla_{j}\log\pi(x)-\nabla_{j}\log\pi^{\prime}(x)\right\|\pi(x)\mathrm{d}x.$ (1.2)
Here $[d]:=\{1,\dots,d\}$, and $\delta$ is a constant independent of the ambient dimension $d$ for distributions with locality structures, including the sparse graphical models [29].
The inequality (1.2) can be interpretted as a marginal version of (1.1), where the left hand side concerns the marginals of distributions of interest, and the right hand side concerns the difference in the score’s individual components. Using (1.2), we can obtain a dimension independent uniform control of the marginal errors, which is not achievable with (1.1). This is useful if each marginal corresponds to a different location in spatial data [40] and manifold data [42].
Moreover, the marginal form in (1.2) suggests that to obtain a uniform bound on the marginal error, it suffices to minimize the worst coordinate error of the score function, which inspires the localized score matching in Section 5. In particular, we show in Theorem 5.1 the necessary training data can be dimension independent.
The marginal inequality (1.2) is established using Stein’s method [38, 39], which is a useful mathematical tool analyzing the difference between two distributions. To adapt to the marginal setting, we introduce a marginal Stein equation
$\mathcal{L}_{\pi}u(x):=\Delta u(x)+\nabla\log\pi(x)\cdot\nabla u(x)=\phi(x_{i})-\mathbb{E}_{\pi}[\phi(x_{i})],$ (1.3)
where $\phi$ is a test function that only depends on the $i$-th component $x_{i}$. The solution $u$ can then be used to derive the mean difference of $\phi$ on the $x_{i}$-marginals of $\pi$ and another distribution $\pi^{\prime}$. For example, using Stein’s lemma [39], we have
$\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})]-\mathbb{E}_{\pi}[\phi(x_{i})]=\mathbb{E}_{\pi^{\prime}}[\nabla u\cdot(\nabla\log\pi-\nabla\log\pi^{\prime})].$
In Stein’s method, one considers all $\phi$ in a functional class, and seeks a uniform upper bound for all corresponding solution $u$. For example, we will show in 2.1 that if $\pi$ is a sparse graphical model, for all $1$-Lipchitz functions $\phi$, the solution $u$ satisfies $\sum_{j=1}^{d}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq\delta$, and then we immediately obtain the marginal Otto-Villani inequality
$\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})=\sup_{\left|\phi\right|_{\rm Lip}\leq 1}|\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})]-\mathbb{E}_{\pi}[\phi(x_{i})]|\leq\delta\cdot\max_{j}\mathbb{E}_{\pi^{\prime}}\left[\left\|\nabla_{j}\log\pi-\nabla_{j}\log\pi^{\prime}\right\|\right].$
This motivates the definition of $\delta$-locality, where we use the bound of $u$ to quantify the locality structure. Formally, $u=\mathcal{L}_{\pi}^{-1}(\phi-\mathbb{E}_{\pi}[\phi])$, so that the condition is essentially imposed on $\mathcal{L}_{\pi}^{-1}$, which encodes the locality in $\pi$. A main technical result of this paper is to show that log-concave sparse graphical models are $\delta$-localized with dimension independent $\delta$ (2.1). Thus the $\delta$-locality condition can be regarded a relaxation of the sparse graphical models. We also show that it holds for distributions with certain diagonal dominance property (2.2).
To design efficient approximation and sampling algorithms for localized distributions, a key observation is that the effective dimension of these problems is determined by the sparsity of local dependency rather than the ambient dimension. For instance, by the clique decomposition of the graphical models [4, 29],
$\log\pi(x)=\sum_{j\in[d]}u_{j}(x_{\mathcal{N}_{j}}),$
where $\mathcal{N}_{j}$ is the neighboring indices of $j$. In sparse graphical models, $|\mathcal{N}_{j}|\ll d$, so that the log density is a sum of low-dimensional functions, greatly reducing the complexity of approximating the distribution. Another advantage is that the conditional dependency allow for efficient local and parallel algorithms [41, 17, 19]. In this paper, we discuss two examples of localized sampling methods: the localized likelihood-informed subspace (LIS) method and the localized score matching method, which are exactly the localized versions of LIS methods [11, 10, 12] and score matching [26, 37].
1.1 Related Literature
1.1.1 Locality Structures
Locality structure is an important structure that has been studied accross various fields, including spatial statistics [4, 44, 13], data assimilation [25, 21, 35], and quantum mechanics [28, 3, 6]. In spatial models, it is observed that correlations between parameters diminish significantly at large distances, typically at exponential rate [28, 5, 35]. This allows for computationally efficient approximations by artificially removing long-range dependency.
In spatial statistics, the Vecchia approximation [44] proposes to approximate Gaussian processes by removing conditioners at large distances, i.e.
$\pi(X_{1})\prod_{i=1}^{d-1}\pi(X_{i+1}|X_{1:i})\approx\pi(X_{1})\prod_{i=1}^{d-1}\pi(X_{i+1}|X_{\mathcal{N}_{i+1}}),$
where $\mathcal{N}_{i+1}$ only contains a small subset of preceding points, thus largely reduces the sampling complexity. Various extensions have been proposed based on the nearest-neighbor approximation idea, see the brief review in [13].
In data assimilation, localization methods [25, 21] are introduced to mitigate spurious long-range correlations arising from small ensemble sizes in the ensemble Kalman filter. For instance, the covariance localization artificially remove or dampen long-range correlations in the ensemble covariance $\widehat{C}$, i.e. $\widehat{C}_{\rm loc}=\rho\circ\widehat{C},~{}\rho_{ij}=\psi(|i-j|)$, where $\circ$ denotes the Hadamard product and $\psi:\mathbb{N}\to\mathbb{R}_{+}$ is a rapidly decaying function. Such localization methods have been shown to effectively reduce sampling errors and improve filter accuracy [25, 21, 35].
In quantum mechanics, the nearsightedness principle [28] states the properties of a quantum system are primarily determined by local interactions and are insensitive to distant perturbations. The principle can be formalized as the exponential decay of the density operator $\rho(\bm{r},\bm{r}^{\prime})$ w.r.t. the physical distance $\left\|\bm{r}-\bm{r}^{\prime}\right\|$ [28, 3], which enables efficient sparse approximation of $\rho$. Such approximation can be regarded as the quantum version of localization of distributions, and significantly reduces the computational cost of tasks such as computing the ground states or expectation value of observables in quantum systems. It also lays the theoretical foundation of the interatomic potentials in molecular dynamics simulation [6].
1.1.2 Localized Sampling
In recent years, there has been a fast growing interest in sampling methods that leverage locality structures [46, 32, 31, 41, 19, 17, 8]. [32] proposes to apply the localization technique in Markov chain Monte Carlo (MCMC) to solve Bayesian inverse problems and introduces a localized Metropolis-within-Gibbs sampler. [41] extends this idea and proves a dimension independent convergence rate of the MALA-within-Gibbs for localized distributions, while [31] discusses how to parallelize the proposal computation. Recently, [17] utilizes the localized structure to obtain a dimension independent approximation error for marginals in an image deblurring problem, based on which we develop the idea to the general case in this paper. Meanwhile, [8] demonstrates that when the distribution exhibits localized structure, the bias of the unadjusted Langevin algorithm for the marginals is at most logarithmically dependent on the dimension. Beyond MCMC, [46] proposes Message Passing Stein Variational Gradient Descent (MP-SVGD). By leveraging the conditional independencies, MP-SVGD finds the descent direction coordinate-wisely, and reduces the degeneracy issue of kernel methods in high dimensions. [19] proposes a localized version of the Schrödinger Bridge (SB) sampler. It replaces a single high-dimensional SB problem by $d$ low-dimensional SB problems, avoiding the exponential dependence of the sample complexity on the dimension.
1.1.3 Stein’s Method and Otto-Villani Inequality
Stein’s method is a useful approach for quantifying distances between probability distributions. First developed in [38] for Gaussian approximation, it has been extended to various distributions, including Poisson [7], binomial [39], and high dimensional settings [36, 9]. We refer to Stein’s monograph [39] for a comprehensive review.
Otto and Villani [34] establish a series of inequalities on the relative entropy $\mathsf{H}$, Wasserstein distance $\mathsf{W}_{2}$ and Fisher information $\mathsf{I}$, and reveal their deep connections with the Riemannian-geometric structure of the Wasserstein space of probability measures. The rich structure and wide connections with various fields make itself an important research topic [45]. In this paper, we refer to the Otto-Villani inequality as the “WI” inequality in Definition 3 in [34]:
$\mathsf{W}_{2}(\mu,\nu)\leq C_{\mu}\sqrt{\mathsf{I}(\mu|\nu)},$
where $\mathsf{I}(\mu|\nu)=\mathbb{E}_{\mu}\left\|\nabla\log\frac{\mu}{\nu}\right\|^{2}$ denotes the Fisher information. For more discussions and survey on tranport inequalities, we refer to [45, 20].
1.2 Arrangement and Notations
The paper is organized as follows. We introduce the $\delta$-locality condition in Section 2, and show that it holds for a large class of distributions with locality structure, including the sparse graphical models. In Section 3, we establish the marginal Otto-Villani inequality for $\delta$-localized distributions using the Stein’s method. The next two sections discuss two applications of the theory: localized LIS method in Section 4 and localized score matching in Section 5. In Section 6, we provide the proof of the technical results in Section 2.
Notations
  • Sets. Denote $[d]=\{1,\dots,d\}$. Denote the cardinality of a set $A$ as $|A|$.
  • Norms. For $v\in\mathbb{R}^{d}$, denote the vector $p$-norm as $\left\|v\right\|_{p}$ and briefly denote $\left\|v\right\|=\left\|v\right\|_{2}$. Likewisely, for a matrix $A\in\mathbb{R}^{m\times n}$, $\left\|A\right\|=\sup_{v\neq 0}\frac{\left\|Av\right\|}{\left\|v\right\|}$ denotes the matrix $2$-norm. $\left\|A\right\|_{\rm F}=\sqrt{\sum_{i,j}A_{ij}^{2}}$ denotes the Frobenius norm.
    For a scalar function $f:\mathbb{R}^{d}\to\mathbb{R}$, denote $\left\|f\right\|_{L^{p}(\pi)}$ as the weighted $L^{p}$ norm $\left\|f\right\|_{L^{p}(\pi)}=\left(\int_{\mathbb{R}^{d}}|f(x)|^{p}\pi(x)\mathrm{d}x\right)^{1/p}$. For a vector-valued function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$, similarly denote $\left\|f\right\|_{L^{p}(\pi)}=\left(\int_{\mathbb{R}^{d}}\left\|f(x)\right\|^{p}\pi(x)\mathrm{d}x\right)^{1/p}$. Note here $\left\|f(x)\right\|$ is the vector $2$-norm.
  • Matrix. Denote $A\succeq B$ if $A-B$ is positive semi-definite, and $A\preceq B$ if $B\succeq A$.
  • Probability. Denote $\mathcal{P}_{1}(\mathbb{R}^{d})$ as the set of probability distributions in $\mathbb{R}^{d}$ with finite first moment. The $1$-Wasserstein distance between two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is
    $\mathsf{W}_{1}(\pi,\pi^{\prime})=\inf_{\gamma\in\Pi(\pi,\pi^{\prime})}\int_{\mathbb{R}^{d}}\left\|x-y\right\|\mathrm{d}\gamma(x,y),$
    where $\Pi(\pi,\pi^{\prime})$ is the set of couplings of $\pi$ and $\pi^{\prime}$, i.e. if $(X,Y)\sim\gamma\in\Pi(\pi,\pi^{\prime})$, then $X\sim\pi,Y\sim\pi^{\prime}$.
    Denote conditional independence $X\perp\!\!\!\perp Y~{}|~{}Z\Leftrightarrow p(x,y|z)=p(x|z)p(y|z)$.
2 $\delta$-localized Distributions
In this paper, we consider a high dimensional distribution in $\mathbb{R}^{d}$. In many applications, it is more reasonable to decompose the components into local blocks, and study the dependence between different blocks. Examples include, a small group of genes functioning together as a pathway, and geophysical state of nearby locations representing state of a particular area. To accomodate such settings, we assume there is a fixed block decomposition
$x=(x_{1},\dots,x_{b});\quad\forall i\in[b],~{}x_{i}\in\mathbb{R}^{d_{i}};\quad d=\sum_{i=1}^{b}d_{i},$ (2.1)
where we denote $[b]:=\{1,\dots,b\}$. Denote $-i:=[b]\backslash\{i\}$. For a distribution $\pi$, denote $\pi_{i}$ as the marginal distribution on $x_{i}$, i.e. $\pi_{i}(x_{i})=\int\pi(x_{i},x_{-i})\mathrm{d}x_{-i}$. Denote the partial derivatives $\nabla_{i}f=\nabla_{x_{i}}f\in\mathbb{R}^{d_{i}},\nabla_{ij}^{2}f=\nabla_{x_{i}x_{j}}^{2}f\in\mathbb{R}^{d_{i}\times d_{j}}$.
We begin by defining the $\delta$-locality of a distribution, which quantifies the local parameter dependence of the solution to the marginal Stein equation.
Definition 2.1.
A distribution $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ is called $\delta$-localized for some constant $\delta>0$, independent of $d$, if for any $i\in[b]$ and $1$-Lipschitz function $\phi:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}$, the solution $u(x)$ to the marginal Stein equation
$\Delta u(x)+\nabla\log\pi(x)\cdot\nabla u(x)=\phi(x_{i})-\mathbb{E}_{\pi}[\phi(x_{i})]$ (2.2)
satisfies the gradient estimate
$\left\|\nabla u\right\|_{\infty,1}:=\sum_{j=1}^{b}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq\delta.$ (2.3)
2.1 is a technical description of the locality structure of the distribution motivated by the Stein’s method. We will show in 3.1 that the $\delta$-locality condition directly implies that local perturbation of the distribution only introduces dimension independent marginal error. We will see from the proof of 3.1 that
$\nabla_{j}u(x)$ quantifies how modifications of $\nabla_{j}\log\pi(x)$ affect the marginal $\pi_{i}$.
We keep the technical definition here for the sake of generality. In the following, we consider two classes of distributions, i.e. the sparse graphical models and the distributions with certain diagonal dominance condition, and show that they are $\delta$-localized with $\delta$ independent of $d$.
2.1 Sparse Graphical Models
In this section, we show that the sparse graphical model [29], also known as the Markov random field, is $\delta$-localized. As a widely adopted model for distributions with locality structure, it uses an undirected graph $G=(V,E)$ to characterize the local dependence of a random variable $X=(X_{i})_{i\in V}$, where for any nonadjacent vertices $i,j\in V$, the two components $X_{i},X_{j}$ are conditionally independent given the rest of the components $(X_{k})_{k\neq i,j}$.
Precisely, consider an undirected graph $G=(V,E)$ where $|V|=b$. Attach each vertex in $V$ with a unique index $j\in[b]$. Denote $\mathcal{N}_{j}=\{k\in V:(j,k)\in E\}$ as the neighboring vertices of $j$. For simplicity, we require that $E$ includes all the self-loops in $G$, i.e. $(j,j)\in E~{}(\forall j\in G)$.
We say a distribution $\pi(x)\in\mathcal{P}(\mathbb{R}^{d})$ has dependency graph $G$ if for $X\sim\pi$,
$\forall j,k\in V,~{}k\notin\mathcal{N}_{j}~{}\Rightarrow~{}X_{j}\perp\!\!\!\perp X_{k}~{}|~{}X_{\mathcal{N}_{j}^{-}},$ (2.4)
[…]
$\forall k\notin\mathcal{N}_{j}~{}\Rightarrow~{}\nabla_{jk}^{2}\log\pi(x)=0.$
Another equivalent characterization of the sparse graphical model is the clique decomposition [29] of the log density
$\log\pi(x)=\sum_{j\in V}u_{j}(x_{\mathcal{N}_{j}}).$
The clique decomposition is a key property that allows efficient approximation and computation of the distribution, as it implies that the log density is essentially a low-dimensional function, which will be exploited in Section 5.
2.1.1 The Sparsity Condition
To specify the sparsity condition, denote the $q$-th neighbor $\mathcal{N}_{j}^{q}~{}(q\in\mathbb{Z}_{+})$ as
$\mathcal{N}_{j}^{q}:=\{k\in V:\mathsf{d}_{G}(j,k)\leq q\},$
where $\mathsf{d}_{G}(j,k)$ is the graph path distance between $j,k\in V$, i.e. the minimum number of edges that must be traversed to go from $j$ to $k$. For convenience, define $\mathsf{d}_{G}(j,j)=0$, and $\mathsf{d}_{G}(j,k)=\infty$ if $j,k$ are not connected.
Definition 2.2.
An undirected graph $G$ is called $\bm{(S,\nu)}$-local if
$\forall j\in V,~{}q\in\mathbb{N},\quad|\mathcal{N}_{j}^{q}|\leq 1+Sq^{\nu}.$ (2.5)
In the above definition, $S$ denotes the maximal size of the immediate neighbor, and $\nu\in\mathbb{Z}_{+}$ denotes the ambient dimension of the graph, which controls the growth speed of the neighborhood volume with the radius. Here we require it growing at most polynomially to ensure effective locality. A motivating example for 2.2 is the lattice model $\mathbb{Z}^{\nu}$, where a naive bound of the neighborhood volume is
$|\mathcal{N}_{j}^{q}|=|\{k\in\mathbb{Z}^{\nu}:\left\|k\right\|_{1}\leq q\}|\leq(2q+1)^{\nu}<1+(3q)^{\nu}.$
So that $\mathbb{Z}^{\nu}$ is $(3^{\nu},\nu)$-local. A simple class of distributions with such sparse dependency graph is those with banded Hessian, i.e. for some bandwidth $B$, the Hessian $H(x):=\nabla^{2}\log\pi(x)$ satisfies
$\forall|j-k|>B~{}\Rightarrow~{}H_{jk}(x)\equiv 0.$
In this case, the dependency graph is $(2B,1)$-local.
Example 2.1 (1D Ginzburg-Landau).
The Ginzburg-Landau (GL) model is a widely used model in statistical physics [30]. A discrete 1D GL model describes a chain of real-valued spins $\{x_{j}\}_{j=1}^{n}$, where each $x_{j}\in\mathbb{R}$ interacts locally. Its Gibbs distribution is given by
$\pi(x)=\frac{1}{Z}\exp\left(\sum_{j=1}^{n}V(x_{j})+\sum_{j=1}^{n-1}W(x_{j},x_{j+1})\right),$
where $V(x)=\frac{\lambda}{4}(x^{2}-m^{2})^{2}$ is the double-well potential and $W(x,y)=\frac{\beta}{2}(x-y)^{2}$ is the nearest-neighbor interaction. It is clear that $\log\pi$ has banded Hessian with bandwidth $1$, and thus the distribution is $(2,1)$-local.
2.1.2 Sparse Graphical Models are $\delta$-localized
Theorem 2.1.
Let $G=(V,E)$ be a $(S,\nu)$-local graph. Suppose $\pi\in\mathcal{P}(\mathbb{R}^{d})$ has dependency graph $G$, and satisfies for some $0<m\leq M<\infty$,
$\forall x\in\mathbb{R}^{d},\quad mI\preceq-\nabla^{2}\log\pi(x)\preceq MI.$
Then $\pi$ is $\delta$-localized with $\delta=\frac{S\nu!\kappa^{\nu}}{m}$ where $\kappa=\frac{M}{m}$.
Proofs can be found in Section 6.2.
Remark 2.1.
The condition number $\kappa$ plays a crucial role in localization:
  • $\kappa$ is known to be crucial to preserve the band structure in matrix inversion [14, 5]. In probability language, consider a Gaussian distribution $\mathsf{N}(0,C)$, then a moderate $\kappa$ ensures the equivalence of local correlation ($C$ is nearly banded) and conditional dependencies ($C^{-1}$ is nearly banded).
  • To achieve a dimension-independent $\delta$, $\kappa$ must remain bounded as the dimension $d$ grows. This is true in many spatial models where the high dimensionality comes from an extended system rather than finer discretization.
2.2 Diagonal Dominant Distributions
In this section, we consider distributions whose dependency graphs are not necessarily sparse, but instead have a certain diagonal dominant property. Such condition implies that any block is mostly correlated with itself, leading to the localized feature of the distribution. This condition is already studied in [17] motivated by the image deblurring problem, and here we show that it also implies the $\delta$-locality.
Theorem 2.2.
Consider $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Suppose $H(x):=-\nabla^{2}\log\pi(x)$ is $c$-unifomly diagonal block dominant, i.e. there exists a matrix $M\in\mathbb{R}_{+}^{b\times b}$ s.t. $\forall i,j\in[b],i\neq j$,
$H_{ii}(x)\succeq M_{ii}I_{d_{i}},\quad\|H_{ij}(x)\|\leq M_{ij};\quad\sum_{j\neq i}M_{ij}+c\leq M_{ii},$
where $H_{ij}(x)$ denotes the $(i,j)$-th subblock of $H(x)$. Then $\pi$ is $\delta$-localized with $\delta=c^{-1}$.
Proof can be found in Section 6.3.
Remark 2.2.
The conditions above imply that $\pi$ is log-concave. Denote $M_{ij}^{\prime}:=2M_{ii}\mathbf{1}_{i=j}-M_{ij}$, then $M^{\prime}$ is $c$-diagonal dominant. Gerŝgorin discs theorem [24] implies that the smallest eigenvalue of $M^{\prime}$ is lower bounded by $c$, and thus
$u^{\rm T}H(x)u=~{}$ $\sum_{i,j}u_{i}^{\rm T}H_{ij}(x)u_{j}\geq\sum_{i}M_{ii}\left\|u_{i}\right\|^{2}-\sum_{i\neq j}M_{ij}\left\|u_{i}\right\|\left\|u_{j}\right\|$
[…]
Therefore, $\pi$ is $c$-strongly log-concave.
3 Marginal Otto-Villani Inequality
3.1 Marginal Error of One Block
Theorem 3.1.
Consider two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Assume $\pi^{\prime}$ is $\delta$-localized, then the marginal $\mathsf{W}_{1}$ distance of $\pi,\pi^{\prime}$ satisfies
$\max_{i\in[b]}\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})\leq\delta\cdot\max_{j\in[b]}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)},$ (3.1)
where $\pi_{i}$ and $\pi_{i}^{\prime}$ denote the marginals of $\pi$ and $\pi^{\prime}$ on $x_{i}$ respectively.
Proof.
By Kantorovich duality [45],
$\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})=\sup_{\phi\in{\rm Lip}_{1}}\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i},$
where ${\rm Lip}_{1}$ denotes the $1$-Lipchitz function class. Fix any $\phi\in{\rm Lip}_{1}$. Note
$\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}=~{}$ $\int\left(\phi(x_{i})-\mathbb{E}_{\pi_{i}^{\prime}}[\phi(x_{i})]\right)\pi_{i}(x_{i})\mathrm{d}x_{i}$
[…]
Here $\mathbb{E}_{\pi_{i}^{\prime}}[\phi(x_{i})]=\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})]$ is a constant. Let $u(x)$ solve the marginal Stein equation
$\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)=\phi(x_{i})-\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})].$
By integration by parts,
$\int\left(\phi(x_{i})-\mathbb{E}_{\pi^{\prime}}[\phi(x_{i})]\right)\pi(x)\mathrm{d}x=~{}$ $\int\left(\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)\right)\pi(x)\mathrm{d}x$
[…]
By assumption, the solution satisfies the gradient estimate (2.3). Therefore,
$\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}\leq~{}$ $\sum_{j=1}^{b}\left\|\nabla_{j}u\right\|_{L^{\infty}}\int\left\|\nabla_{j}\log\pi^{\prime}(x)-\nabla_{j}\log\pi(x)\right\|\pi(x)\mathrm{d}x$
[…]
Therefore, we obtain
$\max_{i}\mathsf{W}_{1}(\pi_{i},\pi_{i}^{\prime})\leq~{}$ $\max_{i}\max_{\phi\in{\rm Lip}_{1}}\int\phi(x_{i})\left(\pi_{i}(x_{i})-\pi_{i}^{\prime}(x_{i})\right)\mathrm{d}x_{i}$
[…]
This completes the proof. ∎
Remark 3.1.
To control the marginal error, one can directly apply Otto-Villani inequality [34] on the marginals and obtain
$\mathsf{W}_{2}(\pi_{i},\pi_{i}^{\prime})\leq C_{\pi_{i}}\sqrt{\mathbb{E}_{\pi_{i}}\left\|\nabla\log\pi_{i}-\nabla\log\pi_{i}^{\prime}\right\|^{2}}.$
The main issue of this approach is that we often have access only to $\pi$ and $\pi^{\prime}$ but not to their marginals, and evaluating the marginals is often computationally very challenging as it involves integrating out the other components. This makes the inequality less useful in practice.
3.2 Marginal Error of Multiple Blocks
In 3.1, we show that the marginal distance of $\pi^{\prime}$ and a $\delta$-localized distribution $\pi$ on one block $x_{i}$ is dimension independent. Here we extend this result to the case of multiple blocks.
Theorem 3.2.
Consider two distributions $\pi,\pi^{\prime}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Assume $\pi^{\prime}$ satisfies any one of the conditions in 2.1 or 2.2, then for any index set $I\subset[b]$, the $\mathsf{W}_{1}$ distance of $\pi,\pi^{\prime}$ on the marginal $x_{I}:=(x_{i})_{i\in I}$ satisfies
$\mathsf{W}_{1}(\pi_{I},\pi_{I}^{\prime})\leq\delta|I|\cdot\max_{j\in[b]}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.$ (3.2)
Here $\delta$ can be taken as the same $\delta$ in 2.1 or 2.2.
Proofs can be found in Section 6.4. 3.2 provides further control on the correlation between different blocks in $\pi^{\prime}$, which cannot be directly derived from 3.1. Roughly speaking, when $\pi$ and $\pi^{\prime}$ are both Gaussians, 3.1 only guarantees that the diagonal blocks of the covariance matrix of $\pi^{\prime}$ are close to those of $\pi$, while 3.2 further guarantees that the off-diagonal blocks are also close.
4 Application I: Localized LIS Method
4.1 Review of LIS Method
Likelihood-informed subspace (LIS) method [11, 10, 12] is a dimension reduction technique to sample efficiently from a high-dimensional distribution of the form
$\pi(x)=\frac{1}{Z}l(x)\pi_{0}(x),$ (4.1)
where $\pi_{0}(x)$ is the prior distribution, $l(x)$ is the likelihood function, and $Z$ is the normalizing constant. Sampling from such distributions is a central problem in Bayesian inference.
The key idea in LIS is that the likelihood function $l(x)$ often exhibits a low-rank structure, making it possible to be approximated by a rigde function
$l(x)\approx l_{r}(x_{r}),\quad x_{r}=P_{r}x,$
where $P_{r}$ is a projection to a low-dimensional likelihood-informed subspace $\mathcal{X}_{r}\subset\mathcal{X}$. This leads to an approximation of the target distribution as
$\pi(x)\approx\pi_{r}(x)\propto l_{r}(x_{r})\pi_{0}(x).$
With a well-chosen subspace $\mathcal{X}_{r}$ and ridge approximation $l_{r}(x)$, LIS provides a certified approximation [12], significantly reducing the sampling complexity. The method has been successfully applied to many high-dimensional Bayesian inference problems [11, 10].
4.2 Localized LIS Method
In LIS, the informed subspace is constructed globally, which may not capture the spatial structure in the target distribution. For localized distributions, a locally defined informed subspace can be a better choice. Based on this observation, we propose to use localized likelihood-informed subspace (LLIS), which incorporates domain decomposition in LIS method. In essence, we decompose the parameters into small blocks, and apply LIS in each block to construct the LLIS.
In the following, we show how to apply LLIS when the prior is a Gaussian distribution, and establish a dimension independent error for LLIS approximation under the sparse graphical assumption.
4.2.1 Construction of Localized LIS
Suppose $\pi_{0}=\mathcal{N}(0,C)$, and $\pi_{0},\pi$ both have the dependency graph $G=(V,E)$. We will apply the LLIS approximation under the whitened transformation, i.e. consider the change of variables
$\tilde{x}=C^{-1/2}x.$
Note $\tilde{x}\sim\tilde{\pi}_{0}=\mathcal{N}(0,I_{d})$ when $x\sim\pi_{0}$. Under the whitened coordinate $\tilde{x}$, we aim to construct the LLIS of the product form
$\widetilde{\mathcal{X}}_{r}=\bigoplus_{j\in V}\widetilde{\mathcal{X}}_{r,j},$
where $\widetilde{\mathcal{X}}_{r,j}\subset\mathbb{R}^{d_{j}}$ denotes the subspace of the reduced paramater $\tilde{x}_{r,j}$. $\widetilde{\mathcal{X}}_{r,j}$ is determined via the following diagnostic matrix
~$\begin{split}\widetilde{\mathcal{M}}=~{}&\mathbb{E}_{\tilde{x}\sim\tilde{\pi}}\left[\tilde{\nabla}\log\tilde{l}(\tilde{x})(\tilde{\nabla}\log\tilde{l}(\tilde{x}))^{\rm T}+\tilde{\nabla}^{2}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}^{2}\log\tilde{l}(\tilde{x})\right]\\ =~{}&C^{1/2}\cdot\mathbb{E}_{x\sim\pi}\left[\nabla\log l(x)(\nabla\log l(x))^{\rm T}+\nabla^{2}\log l(x)\cdot C\cdot\nabla^{2}\log l(x)\right]\cdot C^{1/2},\end{split}$ (4.2)
where we denote the transformed quantities with a tilde, to be specific,
$\tilde{x}=C^{-1/2}x,\quad\tilde{\pi}(\tilde{x})=(\det C)^{-\frac{1}{2}}\pi(x),\quad\tilde{l}(\tilde{x})=l(x),\quad\tilde{\nabla}=\nabla_{\tilde{x}},\quad\tilde{\nabla}^{2}=\nabla^{2}_{\tilde{x}\tilde{x}}.$
Notice compared to the existing LIS methods [12], where only the first order information is used, here we also incorporates the second order information in the diagnostic matrix.
With the diagnostic matrix, we define $\widetilde{\mathcal{X}}_{r,j}$ to be the rank-$r_{j}$ principle subspace of ~$\widetilde{\mathcal{M}}_{jj}\in\mathbb{R}^{d_{j}\times d_{j}}$, the diagonal block of $\widetilde{\mathcal{M}}$ for block $j$. To be specific,
$\widetilde{\mathcal{X}}_{r,j}=\text{span}\{u_{1},\dots,u_{r_{j}}\},$
where $u_{1},\dots,u_{r_{j}}$ are the top $r_{j}$ eigenvectors of ~$\widetilde{\mathcal{M}}_{jj}$, and $r_{j}=\text{\rm rank}(\widetilde{\mathcal{X}}_{r,j})$ is the number of selected informed directions, which can be determined, for instance, by (see [12])
$r_{j}=\min\left\{1\leq r\leq d_{j}:\sum_{k=1}^{\nu}\lambda_{k}(\widetilde{\mathcal{M}}_{jj})\geq(1-\epsilon)\text{\rm tr}[\widetilde{\mathcal{M}}_{jj}]\right\},$
where $\lambda_{k}(\widetilde{\mathcal{M}}_{jj})$ denotes the $k$-th largest eigenvalue of ~$\widetilde{\mathcal{M}}_{jj}$, and $\epsilon\in(0,1)$ is some prescribed threshold. Given $\widetilde{\mathcal{X}}_{r}$, denote the according complementary space
$\widetilde{\mathcal{X}}_{\bot,j}=(\widetilde{\mathcal{X}}_{r,j})^{\bot},\quad\widetilde{\mathcal{X}}_{\bot}=\bigoplus_{j\in V}\widetilde{\mathcal{X}}_{\bot,j}.$
Denote the reduced parameter $\tilde{x}_{r}$, its complement $\tilde{x}_{\bot}$ and their counterparts in the original representation as
$\tilde{x}_{r}=(\tilde{x}_{r,j})_{j\in V},\quad\tilde{x}_{r,j}=\tilde{P}_{r,j}\tilde{x}_{j},\quad x_{r}:=C^{1/2}\tilde{x}_{r}\in\mathcal{X}_{r}:=C^{1/2}\widetilde{\mathcal{X}}_{r}.$
$\tilde{x}_{\bot}=\tilde{x}-\tilde{x}_{r}=(\tilde{x}_{\bot,k})_{j\in V},\quad\tilde{x}_{\bot,k}=\tilde{x}_{j}-\tilde{x}_{r,j}=\tilde{P}_{\bot,k}\tilde{x}_{j},\quad x_{\bot}:=C^{1/2}\tilde{x}_{\bot}\in\mathcal{X}_{\bot}:=C^{1/2}\widetilde{\mathcal{X}}_{\bot}.$
Here $\tilde{P}_{r,j}$ is the orthogonal projector onto $\widetilde{\mathcal{X}}_{r,j}$, and $\tilde{P}_{\bot,k}$ is its orthogonal complement. Note that $x_{r}$ and $x_{\bot}$ are not orthogonal in the standard $\ell_{2}$ sense, but are orthogonal in the $C$-weighted inner product, i.e.,
$x_{r}^{\rm T}C^{-1}x_{\bot}=\tilde{x}_{r}^{\rm T}\tilde{x}_{\bot}=0.$
This implies that $x_{r}$ and $x_{\bot}$ are independent under the prior $\pi_{0}=\mathcal{N}(0,C)$.
4.2.2 Localized LIS Approximation
As in [12], we construct the approximate likelihood $l_{r}(x_{r})$ by averaging $l(x)$ in the logarithm regime over the uninformed subspace $\mathcal{X}_{\bot}$, and approximate the target distribution by
$\begin{split}&\pi_{r}(x)\propto l_{r}(x_{r})\pi_{0}(x),\quad x_{r}=P_{r}x,\\ \text{where}\quad&\log l_{r}(x_{r})=\int\log l(x_{r}+x_{\bot})\pi_{0}(x_{\bot})\mathrm{d}x_{\bot}.\end{split}$ (4.3)
Note $\pi_{0}(x_{\bot}|x_{r})=\pi_{0}(x_{\bot})$ since $x_{\bot}$ and $x_{r}$ are independent under $\pi_{0}$. Also note in the whitened coordinate, the approximation takes the same form,
$\begin{split}\log\tilde{l}_{r}(\tilde{x}_{r}):=~{}&\log l_{r}(C^{1/2}\tilde{x}_{r})=\int\log l(C^{1/2}\tilde{x}_{r}+x_{\bot})\pi_{0}(x_{\bot})\mathrm{d}x_{\bot}\\ =~{}&\int\log\tilde{l}(\tilde{x}_{r}+C^{-1/2}x_{\bot})\pi_{0}(x_{\bot})\mathrm{d}x_{\bot}=\int\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}.\end{split}$ (4.4)
Here we choose to average in the logarithm regime for the sake of simplicity of the theoretical results. In practice, one can also consider other averaging schemes, see [12].
4.3 Approximation Error
We show the marginal approximation error of the LLIS method is dimension independent under the sparse graphical assumption. To be specific, we assume
Assumption 4.1.
The target distribution is of the form (4.1), and
  • $\pi_{0}(x),\pi(x)$ have dependency graph $G$ that is $(S,\nu)$-local.
  • $\pi_{0}(x)=\mathcal{N}(0,C)$, and there exists $0<m\leq M<\infty$ s.t.
    $mI\preceq C^{-1}\preceq MI,\quad mI\preceq-\nabla^{2}\log\pi(x)\preceq MI.$
As in [11, 12], we will bound the approximation error of LLIS by the eigenvalue residues of the modified diagnostic matrices (4.2). For simplicity, denote the eigenvalue residue for a positive semidefinite matrix $C\in\mathbb{R}^{d\times d}$ and a subspace $V_{r}\subset\mathbb{R}^{d}$ as
$\mathcal{R}(C,V_{r}):=\text{\rm tr}\left[P_{\bot}CP_{\bot}\right],$ (4.5)
where $P_{\bot}$ is the orthogonal projector onto $V_{r}^{\bot}$.
Theorem 4.1.
Assume the target distribution $\pi$ satisfies 4.1. Consider its LLIS approximation $\pi_{r}$ (4.3). Then there exists a dimension independent constant $C_{\pi}$ s.t. the marginal error of the LLIS approximation satisfies
$\max_{i}\mathsf{W}_{1}(\pi_{i},\pi_{r,i})\leq C_{\pi}\max_{k\in V}\left(\mathcal{R}(\widetilde{\mathcal{G}}_{k},\widetilde{\mathcal{X}}_{r,k})+\sum_{j\in V}\mathcal{R}(\widetilde{\mathcal{H}}_{j,k},\widetilde{\mathcal{X}}_{r,j})\right)^{1/2},$ (4.6)
where we denote $\mathcal{R}$ as the residue defined in (4.5), and the diagnostic matrices
$\begin{split}\widetilde{\mathcal{G}}_{k}:=~{}&\mathbb{E}_{\tilde{x}\sim\tilde{\pi}_{r}}\left[\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x}))^{\rm T}\right].\\ \widetilde{\mathcal{H}}_{j,k}:=~{}&\mathbb{E}_{\tilde{x}\sim\tilde{\pi}_{r}}\left[\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{k}\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})\right].\end{split}$ (4.7)
Proofs can be found in Appendix B.
Remark 4.1.
(1) We would like to compare the right rand side of (4.6) with the diagnostic matrix (4.2):
  • The expectation in (4.7) is taken w.r.t. the approximate distribution $\tilde{\pi}_{r}$, while in (4.2) it is taken w.r.t. the target distribution $\tilde{\pi}$. So that 4.1 is a posteriori error estimate instead of a priori estimate.
  • (4.2) can be written as
    ~$\widetilde{\mathcal{M}}_{jj}=\mathbb{E}_{\tilde{x}\sim\tilde{\pi}}\left[\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x}))^{\rm T}+\sum_{k\in V}\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{k}\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})\right].$
    For the second term here, the summation is taken on the second index ($k$) in ~$\tilde{\nabla}_{j}\tilde{\nabla}_{k}$, while in (4.6), it is taken on the first index ($j$) in ~$\tilde{\nabla}_{j}\tilde{\nabla}_{k}$.
(2) We would also like to comment that the summation $\sum_{j\in V}$ in (4.6) is still a dimension independent term. Since the second order derivatives ~$\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})$ decay exponentially with $\mathsf{d}_{G}(j,k)$, which is a result of the sparsity graphical assumption.
5 Application II: Localized Score Matching
5.1 Review of Score Matching
Score matching is a variational approach to approximating a probability distribution $p(x)$ by fitting its score function $s(x):=\nabla\log p(x)$. Originally introduced in [26] and later extended in [37], score matching circumvents the need for explicit density estimation by minimizing the Fisher divergence between a candidate distribution and the target distribution.
Given a parametric family of distributions $\mathscr{P}=\{p_{\theta}:\theta\in\Theta\}$, the score matching aims to find the parameter $\theta$ that minimizes the Fisher divergence
$\mathsf{F}(p\|p_{\theta}):=\mathbb{E}_{X\sim p}\left\|s_{\theta}(X)-s(X)\right\|^{2},\quad s_{\theta}(x):=\nabla\log p_{\theta}(x).$
While the Fisher divergence is generally not directly computable, an integration by parts yields a tractable objective function:
$J(\theta):=\mathbb{E}_{X\sim p}\left[2\text{\rm tr}\left(\nabla s_{\theta}(X)\right)+\left\|s_{\theta}(X)\right\|^{2}\right].$
Note $J(\theta)$ only differs with the Fisher divergence $\mathsf{F}(p\|p_{\theta})$ by a constant:
$J(\theta)=\mathsf{F}(p\|p_{\theta})-\mathbb{E}_{X\sim p}\left\|\nabla\log p(X)\right\|^{2}.$
Thus minimizing $J(\theta)$ is equivalent to minimizing the Fisher divergence. In practice, given the data $(X_{1},\dots,X_{N})$ sampled from $p$, $J(\theta)$ is estimated by its empirical version:
$\hat{J}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[2\text{\rm tr}\left(\nabla s_{\theta}(X_{i})\right)+\left\|s_{\theta}(X_{i})\right\|^{2}\right].$
The hypothesis class $\mathscr{P}$ is typically designed so that the score function $s_{\theta}(x)$ and its gradient are computationally tractable, enabling efficient optimization of $\hat{J}(\theta)$ and yielding an approximate distribution $\hat{p}_{\theta}$.
5.2 Localized Score Matching
Despite its simple implementation, score matching in high-dimensional spaces becomes computationally expensive due to the complexity of learning a generic score function. To mitigate this issue, we propose leveraging the locality structure in the target distribution. Specifically, if the distribution can be represented or well-approximated by a sparse graphical model, the score function should exhibit a lower complexity. Exploiting this locality, we introduce the localized score matching method, which significantly reduces computational costs while preserving approximation accuracy.
5.2.1 Clique Decomposition of Sparse Graphical Models
The key motivation of localized score matching is the clique decomposition [4, 29] of the sprase graphical models, which a local factorization of the log density. Consider a distribution $p(x)$ with dependency graph $G=(V,E)$. By Hammersley-Clifford theorem [4], one can factorize
$\log p(x)=\sum_{j\in V}u_{j}(x_{\mathcal{N}_{j}}).$ (5.1)
Here $u_{j}$ is a function of $x_{\mathcal{N}_{j}}\in\mathbb{R}^{d_{\mathcal{N}_{j}}}$, and $d_{\mathcal{N}_{j}}$ denotes the local dimension
$d_{\mathcal{N}_{j}}:=\sum_{k\in\mathcal{N}_{j}}d_{k}.$
Suppose $G$ is $(S,\nu)$-local in the sense of 2.2, then the maximal local dimension satisfies
$d_{\rm loc}:=\max_{j}d_{\mathcal{N}_{j}}\leq(1+s)\max_{j}d_{j}.$ (5.2)
Since $d_{\rm loc}\ll d$ for a sparse graphical model, the complexity of learning $\log p(x)$ is greatly reduced. This motivates us to approximate the target distribution using a hypothesis class $\mathscr{P}_{G}$ that explicitly incorporates the locality structure. Specifically, we define
$\mathscr{P}_{G}=\left\{p_{\theta}(x)=\exp\left(\sum_{j\in V}u_{\theta,j}(x_{\mathcal{N}_{j}})\right):u_{\theta,j}\in\mathcal{U}_{\theta,j}\right\},$
where $\mathcal{U}_{\theta,j}$ is a class of functions in $\mathbb{R}^{d_{\mathcal{N}_{j}}}$ to be specified. Distributions in $\mathscr{P}_{G}$ have dependency graph $G$ by construction. Notice the size of the hypothesis space, measured by its metric entropy [43], scales as $O(d^{\alpha}\exp(Cd_{\rm loc}))$, where $\alpha>0$ and $d_{\rm loc}$ is the maximal local dimensional. This structure effectively mitigates the curse of dimensionality, making learning and optimization more tractable in high-dimensional settings. Here we assume that the $G$ is prefixed based on prior knowledge of the target distribution. One can also learn the dependency graph from data, but we do not consider it here.
5.2.2 Localized Score Matching
By the clique decomposition (5.1), the score function of a MGF can be computed as
$s_{j}(x):=\nabla_{j}\log p(x)=\sum_{k\in\mathcal{N}_{j}}\nabla_{j}u_{k}(x_{\mathcal{N}_{k}}).$
Here $s_{j}(x)$ is the $j$-th component of the score function, and we will refer to it as the local score function. To fit the local score, it suffices to minimize
$\mathbb{E}_{X\sim p}\left\|s_{\theta,j}(x)-s_{j}(x)\right\|^{2},$
where $s_{\theta,j}(x)=\nabla_{j}\log p_{\theta}(x)$ and $p_{\theta}\in\mathscr{P}_{G}$. By integrating by parts, it is equivalent to minimize
$J_{j}(\theta):=\mathbb{E}_{X\sim p}\left[2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(X)\right)+\left\|s_{\theta,j}(X)\right\|^{2}\right],$ (5.3)
To estimate the full score function, one may consider minimizing $\max_{j\in V}J_{j}(\theta)$. However, this is not equivalent to minimizing
$\max_{j\in V}\mathbb{E}_{X\sim p}\left\|s_{\theta,j}(X)-s_{j}(X)\right\|^{2}=\max_{j\in V}\left(J_{j}(\theta)+C_{j}\right),\quad C_{j}:=\mathbb{E}_{X\sim p}\left\|s_{j}(X)\right\|^{2}.$
The discrepancy arises because $C_{j}$ varies with $j$ and directly minimizing $\max_{j}J_{j}(\theta)$ may neglect indices $j$ for which $C_{j}$ is large. Nevertheless, it can be relaxed to a saddle point problem:
$\min_{\theta\in\Theta}J_{\rm loc}(\theta):=\min_{\theta\in\Theta}\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(J_{j}(\theta)+\lambda_{j}\right),$ (5.4)
Here, $\Theta$ denotes the parameter space of the hypothesis class. Though $C_{j}$ is not directly accessible, it can be replaced in practice by a computable lower bound. In practice, given the data $(X_{1},\dots,X_{N})$ sampled from $p$, one computes the empirical loss
$\begin{split}&\hat{J}_{\rm loc}(\theta)=\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(\hat{J}_{j}(\theta)+\lambda_{j}\right),\\ &\hat{J}_{j}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(X_{i})\right)+\left\|s_{\theta,j}(X_{i})\right\|^{2}\right].\end{split}$ (5.5)
Note $\hat{J}_{j}$ can be computed locally, in the sense that the local score $s_{\theta,j}$
$s_{\theta,j}(x)=\sum_{k\in\mathcal{N}_{j}}\nabla_{j}u_{\theta,k}(x_{\mathcal{N}_{k}})$
only concerns a small group of the local potential functions $u_{\theta,k}$. This facilitates efficient optimization of the localized score matching loss.
5.3 Reduction in Sample Complexity
In this section, we aim to show that the localized score matching reduces the sample complexity. To this end, we analyze the convergence rate of the empirical minimizer $\hat{\theta}_{*}$ of the empirical loss (5.5) to the population minimizer $\theta_{*}$ of the true loss (5.4), given by
$\theta_{*}=\mathop{\arg\min}_{\theta\in\Theta}J_{\rm loc}(\theta),\quad\hat{\theta}_{*}=\mathop{\arg\min}_{\theta\in\Theta}\hat{J}_{\rm loc}(\theta).$ (5.6)
Our goal is to quantify this convergence rate in terms of the number of samples $N$ and the full ambient dimension $d$. By enlarging the hypothesis class to include the target distribution, i.e. assuming $p\in\mathscr{P}$ so that $p_{\theta_{*}}=p$, this convergence rate provides an upper bound of the sample complexity of learning the target distribution, quantified by the number of samples required to approximate the target distribution within a given error.
For simplicity, we assume that the target distribution $p(x)$ is supported in the unit hypercube $\mathbb{I}^{d}=[0,1]^{d}$, and consider the hypothesis class
$\begin{split}\mathscr{P}=\Big{\{}p_{\theta}(x)=\exp&\left(\sum_{j\in V}u_{\theta,j}(x_{\mathcal{N}_{j}})\right):mI\preceq-\nabla^{2}\log p_{\theta}\preceq MI,~{}u_{\theta,j}\in\mathcal{U}_{\theta,j}\Big{\}},\\ &\mathcal{U}_{\theta,j}=\left\{u_{\theta,j}\in C^{2}_{0}(\mathbb{I}^{d_{\mathcal{N}_{j}}}):\|u_{\theta,j}\|_{C^{2}}\leq R\right\}.\end{split}$ (5.7)
In practice, this class can be implemented using Input Convex Neural Networks (ICNNs) [1].
Theorem 5.1.
Let $G=(V,E)$ be a $(S,\nu)$-local graph. Consider approximating the target distribution $p(x)\in\mathcal{P}_{1}(\mathbb{I}^{d})$ using localized score matching with the hypothesis class defined in (5.7) and assume $p\in\mathscr{P}$. Then with probability at least $1-\lambda$, it holds that
$\max_{i}\mathsf{W}_{1}(p_{i},\hat{p}_{i})\leq CN^{-1/4}\left(\log\lambda^{-1}+\log b\right)^{1/4},$
where $\hat{p}:=p_{\hat{\theta}_{*}}$ (5.6) is the learned distribution. Here
  • $C$ is a constant depending only on $S,\nu,m,M,R$ and the local dimension $d_{\rm loc}$, but not on the total dimension $d$.
  • $b$ is the number of blocks, and is the only term that scales as $O(d)$.
Proofs can be found in Appendix C. This result highlights a key advantage: despite the high ambient dimension $d$, the convergence rate depends only on the local dimension, making the method scalable in high dimensions. The theorem also implies that the sample complexity of localized score matching is bounded by $N_{\varepsilon}=O(\varepsilon^{-4}\log(d/\lambda))$, with only logarithm dependence on the total dimension $d$. This result precisely characterizes the fact that learning a localized distribution circumvents the curse of dimensionality.
6 Proofs
6.1 Explicit Solution of the Stein Equation
Lemma 6.1.
Under the conditions in 2.1 or 2.2, for any $1$-Lipschitz function $\phi:\mathbb{R}^{d_{i}}\to\mathbb{R}$, the solution of the marginal Stein equation (2.2) is given by (up to a constant)
$u(x)=-\int_{0}^{\infty}\mathbb{E}\left(\phi(X_{t,i}^{x})-\mathbb{E}_{\pi}[\phi(x_{i})]\right)\mathrm{d}t.$ (6.1)
where $X_{t}^{x}$ is the path solution of the overdamped Langevin dynamics
$\mathrm{d}X_{t}^{x}=\nabla\log\pi(X_{t}^{x})\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t},\quad X_{0}^{x}=x.$ (6.2)
As a corollary, the gradient estimate holds
$\left\|\nabla_{j}u(x)\right\|\leq\int_{0}^{\infty}\mathbb{E}\|\nabla_{x_{j}}X_{t,i}^{x}\|\mathrm{d}t.$ (6.3)
Remark 6.1.
Note the generator of the Langevin dynamics (6.2) is $\mathcal{L}_{\pi}=\Delta+\nabla\log\pi\cdot\nabla$. Then the marginal Stein equation (2.2) can be written as
$\mathcal{L}_{\pi}u=\Delta u+\nabla\log\pi\cdot\nabla u=\phi(x_{i})-\mathbb{E}_{\pi}[\phi(x_{i})].$
This type of Stein equation is known to admit explicit solutions of the form (6.1) (see [2]).
Proof.
Let $X_{t}^{x}$ solves the Langevin dynamics (6.2). By Dynkin’s formula [33],
$\begin{split}\mathbb{E}u(X_{T}^{x})-u(x)=~{}&\mathbb{E}\int_{0}^{T}\left(\nabla\log\pi(X_{t}^{x})\cdot\nabla u(X_{t}^{x})+\Delta u(X_{t}^{x})\right)\mathrm{d}t\\ =~{}&\int_{0}^{T}\mathbb{E}\left(\phi(X_{t,i}^{x})-\mathbb{E}_{\pi}[\phi(x_{i})]\right)\mathrm{d}t.\end{split}$ (6.4)
Under the conditions in 2.1 or 2.2, the distribution is strongly log-concave (see 2.2), and it is well-known that Law$(X_{t}^{x})$ converges to the equilibrium $\pi$ exponentially. It implies that the limit $T\rightarrow\infty$ exists for both sides in (6.4), and the limit is
$\int u(x)\pi(x)\mathrm{d}x-u(x)=\int_{0}^{\infty}\mathbb{E}\left(\phi(X_{t,i}^{x})-\mathbb{E}_{\pi}[\phi(x_{i})]\right)\mathrm{d}t.$
This gives (6.1) up to a constant. Taking derivative w.r.t $x_{j}$ gives
$\nabla_{j}u(x)=-\int_{0}^{\infty}\mathbb{E}\left[\nabla_{x_{j}}X_{t,i}^{x}\cdot\nabla\phi(X_{t,i}^{x})\right]\mathrm{d}t.$
Note taking derivative on both sides is valid due to the exponential decay of $\nabla_{x_{j}}X_{t,i}^{x}$. Since $\phi$ is $1$-Lipschitz, we obtain
$\left\|\nabla_{x_{j}}u(x)\right\|\leq\int_{0}^{\infty}\mathbb{E}\left[\|\nabla_{x_{j}}X_{t,i}^{x}\|\left\|\nabla\phi(X_{t,i}^{x})\right\|\right]\mathrm{d}t\leq\int_{0}^{\infty}\mathbb{E}\|\nabla_{x_{j}}X_{t,i}^{x}\|\mathrm{d}t.$
This completes the proof. ∎
6.2 Proof of 2.1
Before the proof, we introduce the following technical lemma, which essentially controls the diffusion speed in the sparse graphs.
Lemma 6.2.
Let $H_{t}\in\mathbb{R}^{d\times d}$ be a time-dependent positive definite matrix satisfying:
  1. a.
    $H_{t}$ has dependency graph $G$, i.e. $H_{t}(i,j)=0$ if $\mathsf{d}_{G}(i,j)>1$.
  2. b.
    $\exists M>0$ s.t. $\forall t\geq 0,0\preceq H_{t}\preceq MI_{d}$.
Consider the matrix ODE
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}=-H_{t}G_{t},\quad G_{0}=I_{d}.$ (6.5)
Then for any $t\geq 0$, it holds that
$\|G_{t}(i,j)\|\leq\exp(-tM)\sum_{k=\mathsf{d}_{G}(i,j)}^{\infty}\frac{t^{k}M^{k}}{k!}.$
Proofs can be found in Section A.1, which is based on the polynomial approximation trick in [14] and Dyson series [15].
Proof of 2.1.
Under the conditions in 2.1, 6.1 holds. By (6.3), it remains to control $\nabla_{x_{j}}X_{t,i}^{x}$. Taking derivative w.r.t. $x$ in (6.2), we get
$\mathrm{d}\nabla X_{t}^{x}=-H_{t}\cdot\nabla X_{t}^{x}\mathrm{d}t,\quad H_{t}:=-\nabla^{2}\log\pi(X_{t}^{x}).$ (6.6)
Note here $\nabla X_{t}^{x}=\nabla_{x}X_{t}^{x}\in\mathbb{R}^{d\times d}$. Denote $G_{t}=\mathrm{e}^{mt}\nabla X_{t}^{x}$ and $\tilde{H}_{t}=H_{t}-mI$, then it holds that
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}=\mathrm{e}^{mt}\left(m\nabla X_{t}^{x}-H_{t}\nabla X_{t}^{x}\right)=-\tilde{H}_{t}G_{t},\quad G_{0}=\nabla X_{0}^{x}=I_{d}.$
By assumption, $0\preceq\tilde{H}_{t}\preceq(M-m)I$, and $\tilde{H}_{t}(i,j)=0$ if $\mathsf{d}_{G}(i,j)>1$. By 6.2,
$\mathrm{e}^{mt}\|\nabla_{x_{j}}X_{t,k}^{x}\|=\|G_{t}(k,j)\|\leq\mathrm{e}^{-(M-m)t}\sum_{q=\mathsf{d}_{G}(j,k)}^{\infty}\frac{t^{q}(M-m)^{q}}{q!}.$
The estimate holds for any initial condition $x$. For different vertices $j\in V$, consider different initial conditions $x^{(j)}$, and take summation over $j\in V$, we obtain
$\sum_{j\in V}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\leq~{}$ $\mathrm{e}^{-mt}\mathrm{e}^{-(M-m)t}\sum_{j\in V}\sum_{q=\mathsf{d}_{G}(j,k)}^{\infty}\frac{t^{q}(M-m)^{q}}{q!}$
[…]
Here we use the sparsity condition $|\mathcal{N}_{j}^{q}\backslash\{v\}|\leq Sq^{\nu}$ (2.5). From (6.3), we obtain
$\sum_{j\in V}\|\nabla_{j}u(x^{(j)})\|\leq~{}$ $\sum_{j\in V}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\mathrm{d}t=\mathbb{E}\int_{0}^{\infty}\sum_{j\in V}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\mathrm{d}t$
[…]
By A.2, it holds that (denote $\kappa=\frac{M}{m}$)
$\sum_{j\in V}\|\nabla_{j}u(x^{(j)})\|\leq~{}$ $\frac{1}{m}+\frac{S}{M}\nu!\left(\frac{m}{M}\right)^{-\nu-1}\left(1-\frac{m}{M}\right)$
[…]
Taking supremum over $x^{(j)}$, we obtain the desired gradient estimate with $\delta=\frac{S\nu!\kappa^{\nu}}{m}$. ∎
6.3 Proof of 2.2
Proof.
Similar as the proof of 2.1, it remains to control $\nabla_{x_{j}}X_{t,k}^{x}$ from (6.6). To control the $2$-norm of $\nabla_{x_{j}}X_{t,k}^{x}\in\mathbb{R}^{d_{k}\times d_{j}}~{}(k=1,\dots b)$, consider fixing a test vector $v_{j}\in\mathbb{R}^{d_{j}}$ s.t. $\left\|v_{j}\right\|=1$, and denote $g_{t}^{x}(k,j)=\nabla_{x_{j}}X_{t,k}^{x}\cdot v_{j}\in\mathbb{R}^{d_{k}}$, then
$\frac{\mathrm{d}}{\mathrm{d}t}g_{t}^{x}(k,j)=\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{x_{j}}X_{t,k}^{x}v_{j}=-H_{t}\cdot\nabla_{x_{j}}X_{t}^{x}v_{j}=-\sum_{l=1}^{b}H_{t}(k,l)\nabla_{x_{j}}X_{t,l}^{x}v_{j}=-\sum_{l=1}^{b}H_{t}(k,l)g_{t}^{x}(l,j).$
Here we denote $H_{t}(k,l)\in\mathbb{R}^{d_{k}\times d_{l}}$ as the $(k,l)$-th subblock of $H_{t}$. Then by assumption,
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left\|g_{t}^{x}(k,j)\right\|^{2}=~{}$ $-\sum_{l=1}^{b}(g_{t}^{x}(k,j))^{\rm T}H_{t}(k,l)g_{t}^{x}(l,j)$
[…]
Notice $\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left\|g_{t}^{x}(k,j)\right\|^{2}=\left\|g_{t}^{x}(k,j)\right\|\frac{\mathrm{d}}{\mathrm{d}t}\left\|g_{t}^{x}(k,j)\right\|$, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|g_{t}^{x}(k,j)\right\|\leq-M_{kk}\left\|g_{t}^{x}(k,j)\right\|+\sum_{l\neq k}M_{kl}\left\|g_{t}^{x}(l,j)\right\|.$
Such inequality holds for any index $j,k\in[b]$, test vector $v_{j}$ and initial condition $x$. For different indices $j\in[b]$, consider different initial conditions $x^{(j)}$, and denote the matrix $G_{t}\in\mathbb{R}^{b\times b}$ where
$G_{t}(k,j)=\|g_{t}^{x^{(j)}}(k,j)\|.$
Then the above inequality can be written compactly in a matrix form
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}\leq-\tilde{M}G_{t},\quad\tilde{M}_{ij}:=2M_{ii}\mathbf{1}_{i=j}-M_{ij}.$
Here $\leq$ is in the entrywise sense. Note the initial condition is $G_{0}=I_{b}$, since
$G_{0}(k,j)=\|g_{0}^{x^{(j)}}(k,j)\|=\|\nabla_{x_{j}^{(j)}}X_{0,k}^{x^{(j)}}v_{j}\|=\|\nabla_{x_{j}^{(j)}}x_{k}^{(j)}v_{j}\|=\delta_{jk}\left\|v_{j}\right\|=\delta_{jk}.$
By assumption, $\forall i,j\in[b],i\neq j$, $\tilde{M}_{ij}=-M_{ij}\leq 0$, and
$\sum_{j\neq i}|\tilde{M}_{ij}|+c=\sum_{j\neq i}M_{ij}+c\leq M_{ii}=\tilde{M}_{ii}.$
Thus we can apply A.3 and obtain
$\|G_{t}\|_{\infty}\leq\mathrm{e}^{-ct}\|G_{0}\|_{\infty}=\mathrm{e}^{-ct}~{}\Rightarrow~{}\max_{k}\sum_{j}|\nabla_{x_{j}^{(j)}}X_{t,k}^{x^{(j)}}v_{j}|\leq\mathrm{e}^{-ct}.$
Since $v_{j}$ is arbitrary, we obtain that
$\max_{k}\sum_{j}\|\nabla_{x_{j}^{(j)}}X_{t,k}^{x^{(j)}}\|\leq\mathrm{e}^{-ct}.$
Recall (6.3), this implies
$\sum_{j}\|\nabla_{j}u(x^{(j)})\|\leq~{}$ $\sum_{j}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\mathrm{d}t$
[…]
Now as $x^{(j)}$ is arbitrary, we obtain the desired gradient estimate with $\delta=c^{-1}$. ∎
6.4 Proof of 3.2
Proof.
The proof is based on the proof of 2.1 and 3.1. And we only represent the different parts for the multiple block case.
For any index set $I\subset[b]$, denote $d_{I}=\sum_{i\in I}d_{i}$. Then by definition,
$\mathsf{W}_{1}(\pi_{I},\pi_{I}^{\prime})=\sup_{\phi\in{\rm Lip}_{1}(\mathbb{R}^{d_{I}})}\int\phi(x_{I})\left(\pi_{I}(x)-\pi_{I}^{\prime}(x)\right)\mathrm{d}x_{I}.$
Given $\phi_{I}$, let $u(x)$ solves the marginal Stein equation
$\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)=\phi(x_{I})-\mathbb{E}_{\pi^{\prime}}[\phi(x_{I})].$
Similarly as in 6.1, it holds that
$\left\|\nabla_{j}u(x)\right\|\leq\int_{0}^{\infty}\mathbb{E}\left[\|\nabla_{j}X_{t,I}^{x}\|\left\|\nabla\phi(X_{t,I}^{x})\right\|\right]\mathrm{d}t\leq\int_{0}^{\infty}\mathbb{E}\|\nabla_{j}X_{t,I}^{x}\|\mathrm{d}t.$
where $X_{t}^{x}$ is the path solution of the Langevin dynamics for $\pi^{\prime}$ with initial condition $x$. Notice $\|\nabla_{j}X_{t,I}^{x}\|\leq\sum_{i\in I}\|\nabla_{j}X_{t,i}^{x}\|$, we obtain that
$\left\|\nabla u\right\|_{\infty,1}=\sum_{j\in[b]}\left\|\nabla_{j}u(x)\right\|\leq\sum_{j\in[b]}\mathbb{E}\int_{0}^{\infty}\sum_{i\in I}\|\nabla_{j}X_{t,i}^{x}\|\mathrm{d}t\leq\sum_{i\in I}\delta=\delta|I|.$
Using the same arguments in 3.1, we obtain
$\mathsf{W}_{1}(\pi_{I},\pi_{I}^{\prime})\leq~{}$ $\left\|\nabla u\right\|_{\infty,1}\cdot\max_{j}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}$
[…]
This completes the proof. ∎
7 Conclusion
In this paper, we consider the locality structure in high-dimensional distributions. We introduce a quantitative analysis by proposing the $\delta$-locality condition and establishing a marginal Otto-Villani inequality. We develop the Stein’s method for analyzing the marginals, and give new characterization of the locality structure based on the Stein’s method. The marginal form of the Otto-Villani inequality provides refined approximation guarantees for low-dimensional marginals, which lays the theoretical foundation for localization methods. We also propose to apply the localization methodology in distribution approximation and sampling to design scalable algorithms in high dimensions. As examples, we show how to localize the LIS method and score matching, to design scalable algorithms for high-dimensional distributions.
It wound be interesting to extent the analysis framework to more general distributions and different locality conditions. Relaxing the log concavity assumption, and finding operations that preserve the locality would be practically interested. It is also interesting to investigate the localization methodology in other applications, such as the kernel methods and the modern generative models. We will leave these as future work.
Acknowledgements
The work of SL is partially supported by NUS Overseas Research Immersion Award (ORIA).
Appendix A Proof of Lemmas
A.1 Lemmas for 2.1
Proof of 6.2.
I. Scaling. First note by a scaling argument, it suffices to consider $t=1$. Suppose we aim to obtain the control at $t=t_{0}$. Let $G_{t}$ solves (6.5), then $G_{t_{0}s}$ solves
$\frac{\mathrm{d}}{\mathrm{d}s}G_{t_{0}s}=-t_{0}H_{t_{0}s}G_{t_{0}s},\quad G_{0}=I_{d}.$
If the theorem holds for $t=1$, then we obtain that at $s=1$ (note $\left\|t_{0}H_{t_{0}s}\right\|\leq Mt_{0}$),
$\|G_{t_{0}}(i,j)\|\leq\exp(-Mt_{0})\sum_{k=\mathsf{d}_{G}(i,j)}^{\infty}\frac{M^{k}t_{0}^{k}}{k!}.$
II. Dyson series solution. By variation of constants, we have
$G_{t}=I_{d}-\int_{0}^{t}H_{s}G_{s}\mathrm{d}s.$
Applying this identity recursively, we obtain
$G_{t}=~{}$ $I_{d}-\int_{0}^{t}H_{s}\left(I_{d}-\int_{0}^{s}H_{u}G_{u}\mathrm{d}u\right)\mathrm{d}s$
[…]
$+(-1)^{N}\int_{[0,t]^{N}}H_{t_{N}}\cdots H_{t_{1}}G_{t_{1}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{N}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{N}.$
For simplicity, denote
$X_{0}(t)=I_{d},\quad X_{n}(t):=\frac{n!}{t^{n}}\int_{[0,t]^{n}}H_{t_{n}}\cdots H_{t_{1}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n},\quad n\geq 1.$
$R_{N}(t)=\int_{[0,t]^{N}}H_{t_{N}}\cdots H_{t_{1}}G_{t_{1}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{N}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{N}.$
Notice $X_{n}(t)$ is the average of “$H^{n}$” in $[0,t]$. Then
$G_{t}=\sum_{n=0}^{N-1}(-1)^{n}\frac{t^{n}}{n!}X_{n}(t)+R_{N}(t).$
Now we prove that $\lim_{N\rightarrow\infty}R_{N}(t)=0$. First notice $\left\|G_{t}\right\|_{2}\leq 1$, since
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{\rm T}G_{t}=-2G_{t}^{\rm T}H_{t}G_{t}~{}\Rightarrow~{}G_{t}^{\rm T}G_{t}=I_{d}-\int_{0}^{t}G_{s}^{\rm T}H_{s}G_{s}\mathrm{d}s\preceq I_{d}.$
So that as $N\rightarrow\infty$,
$\left\|R_{N}(t)\right\|_{2}\leq~{}$ $\int_{[0,t]^{N}}\left\|H_{t_{N}}\right\|_{2}\cdots\left\|H_{t_{1}}\right\|_{2}\left\|G_{t_{1}}\right\|_{2}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{N}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}$
[…]
This proves that the Dyson series converges, and
$G_{t}=\sum_{n=0}^{\infty}(-1)^{n}\frac{t^{n}}{n!}X_{n}(t).$
III. Representation of polynomials. Denote the matrix process space
$\mathscr{X}=\text{span}\{X_{n}(t),n\geq 0\}=\overline{\Big{\{}\sum_{k=0}^{n}a_{k}X_{n}(t):a_{k}\in\mathbb{C}\Big{\}}}.$
We define the representation of any polynomial $P$ in $\mathscr{X}$ as
$P[X](t)=\sum_{k=0}^{n}a_{k}X_{k}(t),\quad\text{if }P(x)=\sum_{k=0}^{n}a_{k}x^{k}.$
In A.1 we show that the representation has an equivalent definition:
$P\{X\}(t)=\frac{a_{n}}{t^{n}}\sum_{\sigma\in S_{n}}\int_{0}^{t}(H_{t_{1}}-x_{\sigma_{1}}I)\int_{0}^{t_{1}}(H_{t_{2}}-x_{\sigma_{2}}I)\cdots\int_{0}^{t_{n-1}}(H_{t_{n}}-x_{\sigma_{n}}I)\mathrm{d}t_{n}\cdots\mathrm{d}t_{2}\mathrm{d}t_{1}.$
if $\text{if }P(x)=a_{n}\prod_{k=1}^{n}(x-x_{k}).$
Here $S_{n}$ is the set of all permutations of $[n]$. We use $P\{X\}$ to distinguish different ways of definitions. But they are in fact equivalent. Note also that the representation can be extended from polynomials to any analytic functions.
IV. Banded matrix approximation. By Taylor expansion,
$\exp(-x)=\exp(-M)\sum_{n=0}^{\infty}\frac{(M-x)^{n}}{n!}.$
Represent the series in $\mathscr{X}$, we obtain (denote $\hat{X}_{n}(x)=(M-x)^{n}$)
$G_{1}=\exp(-M)\sum_{n=0}^{\infty}\frac{1}{n!}\hat{X}_{n}[X](1),$
$\hat{X}_{n}[X](1)=\sum_{\sigma\in S_{n}}\int_{[0,1]^{n}}(MI-H_{t_{1}})\cdots(MI-H_{t_{n}})\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}.$
Here we use the alternative representation. We can truncate the Dyson series of $G_{1}$ as
$G_{1}=\exp(-M)\sum_{q=0}^{n}\frac{1}{q!}\hat{X}_{q}[X](1)+\exp(-M)\sum_{q=n+1}^{\infty}\frac{1}{q!}\hat{X}_{q}[X](1).$
Consider the off-diagonal entry $G_{1}(i,j)$. Take $n=\mathsf{d}_{G}(i,j)-1$, then since all the path in $G$ connecting $i$ and $j$ has length no less than $\mathsf{d}_{G}(i,j)>n$, it must hold that
$\forall 1\leq q\leq n,\quad\left[(MI-H_{t_{1}})\cdots(MI-H_{t_{q}})\right](i,j)=0~{}\Rightarrow~{}\hat{X}_{q}[X](1)(i,j)=0,$
Therefore,
$G_{1}(i,j)=\exp(-M)\sum_{q=n+1}^{\infty}\frac{1}{q!}\hat{X}_{q}[X](1)(i,j).$
Since $0\preceq H_{t_{k}}\preceq MI$, it holds that $\left\|MI-H_{t_{k}}\right\|_{\rm op}\leq M$, and thus
$\|\hat{X}_{n}[X](1)\|_{\rm op}\leq M^{n}n!\int_{[0,1]^{n}}\mathbf{1}_{t_{1}\leq t_{2}\leq\cdots\leq t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}=M^{n}.$
$\Rightarrow~{}\|G_{1}(i,j)\|\leq\exp(-M)\sum_{q=n+1}^{\infty}\frac{M^{q}}{q!}=\exp(-M)\sum_{q=\mathsf{d}_{G}(i,j)}^{\infty}\frac{M^{q}}{q!}.$
This verifies the case $i\neq j$. For $i=j$, the result follows from $\|G_{1}(i,i)\|\leq\left\|G_{1}\right\|\leq 1$. ∎
Lemma A.1.
The two representations of polynomials in $\mathscr{X}$ are equivalent, i.e.
$P[X](t)=P\{X\}(t),\quad\forall P\in\mathbb{P}.$
Proof.
We prove by induction. First note $n=1$ is obvious,
$(a_{1}(x-x_{1}))\{X\}(t)=\frac{a_{1}}{t}\int_{0}^{t}(H_{t_{1}}-x_{1}I)\mathrm{d}t_{1}=a_{1}(X_{1}(t)-x_{1}I_{d})=(a_{1}(x-x_{1}))[X](t).$
Now consider $n\geq 2$. Notice $P[X](t),P\{X\}(t)$ can both be viewed as mulitlinear maps on $x_{1},\dots,x_{n}$. Compute the partial derivative w.r.t. $x_{n}$,
$\nabla_{x_{n}}\left(P\{X\}(t)\right)$
[…]
Here the first equality follows discussion on the position of $x_{n}$. The third equality follows integrating the variable $t_{k}$ and notice the constraint $t_{k+1}\leq t_{k}\leq t_{k-1}$. The forth equality follows relabeling the index and taking $t_{n}=0$. The last equality follows the induction hypothesis. By symmetry, the relation holds for any $x_{i}$:
$\nabla_{x_{i}}\left(P\{X\}(t)\right)=-\Big{(}a_{n}\prod_{j\neq i}(x-x_{j})\Big{)}[X](t)=(\nabla_{x_{i}}P)[X](t),$
since $\nabla_{x_{i}}P(x)=-a_{n}\prod_{j\neq i}(x-x_{j})$. Now notice the representation $P[X]$ is linear in the coefficients of $P$, it is direct to verify
$\nabla_{x_{i}}(P[X](t))=(\nabla_{x_{i}}P)[X](t)=\nabla_{x_{i}}\left(P\{X\}(t)\right).$
Finally notice when $x_{1}=\cdots=x_{n}=0$,
$x^{n}\{X\}(t)=\frac{n!}{t^{n}}\int_{0}^{t}H_{t_{1}}\int_{0}^{t_{1}}H_{t_{2}}\cdots\int_{0}^{t_{n-1}}H_{t_{n}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}=X_{n}(t)=x^{n}\{X\}(t).$
Now the two multi-linear maps agree on one point, and also all the partial derivatives. So that they must be identical. This shows the equivalence holds for $n$, and by induction, it holds for all $n$. ∎
Lemma A.2.
For any $t\geq 0$ and $x\in(0,1)$, it holds that
$\sum_{k\geq 1}k^{t}(1-x)^{k}<2\Gamma(t+1)x^{-t-1}(1-x).$
When $t\in\mathbb{N}$, the factor $2$ can be omitted, i.e. $\sum_{k\geq 1}k^{t}(1-x)^{k}\leq t!x^{-t-1}(1-x)$.
Proof.
We prove by induction. For $t=0$, it holds that
$\sum_{k\geq 1}(1-x)^{k}=\frac{1-x}{x}.$
For $t\in(0,1)$, first notice by Abel transformation,
$\sum_{k\geq 1}k^{t}(1-x)^{k}=x^{-1}\sum_{k\geq 1}k^{t}\left[(1-x)^{k}-(1-x)^{k+1}\right]=x^{-1}\sum_{k\geq 1}\left(k^{t}-(k-1)^{t}\right)(1-x)^{k}.$
Therefore, since $k^{t}-(k-1)^{t}\leq t(k-1)^{t-1}$ when $t\in(0,1)$ and $k\geq 2$,
$\sum_{k\geq 1}k^{t}(1-x)^{k}\leq~{}$ $x^{-1}\left[(1-x)+\sum_{k\geq 2}t(k-1)^{t-1}(1-x)^{k}\right]$
[…]
Here we use $t\Gamma(t)=\Gamma(t+1)$ and $\Gamma(t+1)x^{-t}>1$ in the last step. This verifies the case $t\in[0,1)$. Suppose the inequality holds for $t-1\geq 0$, then using the same methods,
$\sum_{k\geq 1}k^{t}(1-x)^{k}=~{}$ $x^{-1}\sum_{k\geq 1}\left(k^{t}-(k-1)^{t}\right)(1-x)^{k}\leq x^{-1}\sum_{k\geq 1}tk^{t-1}(1-x)^{k}$
[…]
Here the first inequality follows by the elementary inequality $k^{t}-(k-1)^{t}\leq tk^{t-1}$ when $t\geq 1$, and the second inequality follows by induction hypothesis. The finer inequality for $t\in\mathbb{N}$ can be obtained similarly. This completes the proof. ∎
A.2 Lemma for 2.2
Lemma A.3.
Suppose $G_{t}\in\mathbb{R}_{+}^{b\times b}$ is a time-dependent nonnegative matrix satisfying
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}\leq-MG_{t},$
where $\leq$ is in the entrywise sense, and $M\in\mathbb{R}^{b\times b}$ is $c$-diagonal dominant with non-positive off-diagonal entries, i.e. $\forall i,j\in[b],i\neq j$,
$\sum_{j\neq i}|M_{ij}|+c\leq M_{ii};\quad M_{ij}\leq 0.$
Then for any $t\geq 0$, it holds
$\|G_{t}\|_{\infty}\leq\mathrm{e}^{-ct}\|G_{0}\|_{\infty}.$
Proof.
Denote $G_{t}^{c}:=\mathrm{e}^{ct}G_{t}$, then
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{c}=\mathrm{e}^{ct}\left(\frac{\mathrm{d}}{\mathrm{d}t}G_{t}+cG_{t}\right)\leq\mathrm{e}^{ct}\left(-MG_{t}+cG_{t}\right)=\left(-M+cI\right)G_{t}^{c}.$
Multiple both sides by $\mathbf{1}=(1,\dots,1)^{\rm T}\in\mathbb{R}^{b}$ from right,
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{c}\mathbf{1}\leq\left(-M+cI\right)G_{t}^{c}\mathbf{1}.$
This operator preserves the inequality since it is equivalent to taking summation over row indices. We claim that
$\forall t\geq 0,\quad\|G_{t}^{c}\mathbf{1}\|_{\infty}\leq\|G^{c}_{0}\mathbf{1}\|_{\infty},$ (A.1)
which is a reformulation of the conclusion $\left\|G_{t}\right\|_{\infty}\leq\left\|G_{0}\right\|_{\infty}$ by noticing $G_{t}^{c}\in\mathbb{R}_{+}^{b}$ and
$\mathrm{e}^{ct}\|G_{t}\|_{\infty}=\|G_{t}^{c}\|_{\infty}=\max_{i}\sum_{j}G_{t}^{c}(i,j)=\|G_{t}^{c}\mathbf{1}\|_{\infty}.$
We prove (A.1) by contraction. Suppose (A.1) is false, then $\exists s\geq 0$ and $i$ s.t.
$\frac{\mathrm{d}}{\mathrm{d}t}(G^{c}_{s}\mathbf{1})_{i}>0;\quad(G^{c}_{s}\mathbf{1})_{i}=\|G^{c}_{s}\mathbf{1}\|_{\infty}.$
On the other hand, notice by assumption on $M$,
$\frac{\mathrm{d}}{\mathrm{d}t}(G^{c}_{s}\mathbf{1})_{i}\leq~{}$ $\left(-M_{ii}+c\right)(G^{c}_{s}\mathbf{1})_{i}+\sum_{j\neq i}(-M_{ij})(G^{c}_{s}\mathbf{1})_{j}$
[…]
Contradiction. This proves our result. ∎
Appendix B Proof of 4.1
Proof.
By 2.1, $\pi$ is $\delta$-localized with $\delta=\frac{S\nu!\kappa^{\nu}}{m}$. 3.1 then implies
$\max_{i\in V}\mathsf{W}_{1}(\pi_{i},\pi_{r,i})\leq\delta\cdot\max_{j\in V}\left\|\nabla_{j}\log\pi-\nabla_{j}\log\pi_{r}\right\|_{L^{1}(\pi_{r})}.$
Denote $x:=C^{-1/2}\tilde{x}$, $\nabla_{j}=\nabla_{x_{j}}$ and ~$\tilde{\nabla}_{j}=\nabla_{\tilde{x}_{j}}$. By the chain rule,
$\nabla_{j}=\sum_{k\in V}\frac{\partial\tilde{x}_{k}}{\partial x_{j}}\cdot\tilde{\nabla}_{k}=\sum_{k\in V}C^{1/2}(j,k)\tilde{\nabla}_{k}.$
By B.1 (see below), $\forall j,\sum_{k\in V}\|C^{1/2}(j,k)\|\leq m^{-1/2}S\nu!\kappa^{\nu}$. So that
$\max_{j}\int\|\nabla_{j}\log\pi(x)-\nabla_{j}\log\pi_{r}(x)\|\pi_{r}(x)\mathrm{d}x$
[…]
Denote $C_{\pi}=m^{-3/2}(S\nu!\kappa^{\nu})^{2}$, then
$\max_{j}\mathsf{W}_{1}(\pi_{j},\pi_{r,j})\leq C_{\pi}\max_{k}\int\|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})\|\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}.$ (B.1)
Next we control the right hand side of (B.1). Note $\tilde{\pi}\propto\tilde{l}\tilde{\pi}_{0}$ and $\tilde{\pi}_{r}\propto\tilde{l}_{r}\tilde{\pi}_{0}$, it holds that
~$\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})=\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{l}_{r}(\tilde{x}_{r}),$
Due to the orthogonality of $\tilde{x}_{r}$ and $\tilde{x}_{\bot}$, one can decompose
~$\|\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2}=\|\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})\|^{2}+\|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2},$
where we denote ~$\tilde{\nabla}_{r,k}=\nabla_{\tilde{x}_{r,k}},\tilde{\nabla}_{\bot,k}=\nabla_{\tilde{x}_{\bot,k}}$ and use ~$\tilde{\nabla}_{\bot,k}\log\tilde{l}_{r}(\tilde{x}_{r})=0$ since $\tilde{l}_{r}(\tilde{x}_{r})$ does not depend on $\tilde{x}_{\bot}$. Therefore,
~$\begin{split}\int\|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})\|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}=\int\|\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})\|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}&\\ +\int\|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}=:~{}&\mathsf{I}+\mathsf{II}.\end{split}$ (B.2)
The first term is straightforward to control. Notice
$\begin{split}\mathsf{I}=~{}&\int\text{\rm tr}\left[\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{\bot,k}\log\tilde{l}(\tilde{x}))^{\rm T}\right]\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\\ =~{}&\text{\rm tr}\left[\int\tilde{P}_{\bot,k}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})(\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x}))^{\rm T}\tilde{P}_{\bot,k}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\right]\\ =~{}&\text{\rm tr}\left[\tilde{P}_{\bot,k}\widetilde{\mathcal{G}}_{k}\tilde{P}_{\bot,k}\right]=\mathcal{R}(\widetilde{\mathcal{G}}_{k},\widetilde{\mathcal{X}}_{r,k}).\end{split}$ (B.3)
For the second term, first notice $\tilde{\pi}_{r}$ can be decomposed as
$\tilde{\pi}_{r}(\tilde{x})=\tilde{Z}_{r}^{-1}\tilde{l}_{r}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x})=\tilde{Z}_{r}^{-1}\tilde{l}_{r}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x}_{\bot})=\tilde{\pi}_{r}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x}_{\bot}),$
where $\tilde{\pi}_{r}(\tilde{x}_{r})=\tilde{Z}_{r}^{-1}\tilde{l}_{r}(\tilde{x}_{r})\tilde{\pi}_{0}(\tilde{x}_{r})$ denotes the marginal distribution of $\tilde{\pi}_{r}$ on $\tilde{x}_{r}$. So that
$\mathsf{II}=\int\left(\int\|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}\right)\tilde{\pi}_{r}(\tilde{x}_{r})\mathrm{d}\tilde{x}_{r}.$
Take ~$\tilde{\nabla}_{r,k}$ in (4.4), we have
~$\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})=\int\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}.$
This implies that ~$\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})$ is the average of ~$\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})$ with respect to $\tilde{\pi}_{0}(\tilde{x}_{\bot})$. Therefore, we can apply the Poincaré inequality for the function ~$\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})$ with reference measure $\tilde{\pi}_{0}(\tilde{x}_{\bot})$ (which is a standard Gaussian) to obtain
~$\int\|\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x}_{r}+\tilde{x}_{\bot})-\tilde{\nabla}_{r,k}\log\tilde{l}_{r}(\tilde{x}_{r})\|^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}$
[…]
Note one can relax
~$\|\tilde{\nabla}_{\bot,j}\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})\|_{\rm F}^{2}=~{}$ $\text{\rm tr}\left[\tilde{\nabla}_{\bot,j}\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{r,k}\tilde{\nabla}_{\bot,j}\log\tilde{l}(\tilde{x})\right]$
[…]
Therefore,
$\begin{split}\mathsf{II}\leq~{}&\int\left(\int\sum_{j\in V}\|\tilde{\nabla}_{\bot,j}\tilde{\nabla}_{r,k}\log\tilde{l}(\tilde{x})\|_{\rm F}^{2}\tilde{\pi}_{0}(\tilde{x}_{\bot})\mathrm{d}\tilde{x}_{\bot}\right)\tilde{\pi}_{r}(\tilde{x}_{r})\mathrm{d}\tilde{x}_{r}\\ \leq~{}&\int\sum_{j\in V}\text{\rm tr}\left[\tilde{P}_{\bot,j}\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})\cdot\tilde{\nabla}_{k}\tilde{\nabla}_{j}\log\tilde{l}(\tilde{x})\tilde{P}_{\bot,j}\right]\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\\ =~{}&\sum_{j\in V}\text{\rm tr}\left[\tilde{P}_{\bot,j}\widetilde{\mathcal{H}}_{j,k}\tilde{P}_{\bot,j}\right]=\sum_{j\in V}\mathcal{R}(\widetilde{\mathcal{H}}_{j,k},\widetilde{\mathcal{X}}_{r,j}).\end{split}$ (B.4)
Combine (B.1)-(B.4), we obtain
$\max_{j}\mathsf{W}_{1}(\pi_{j},\pi_{r,j})\leq~{}$ $C_{\pi}\max_{k}\left(\int\|\tilde{\nabla}_{k}\log\tilde{\pi}(\tilde{x})-\tilde{\nabla}_{k}\log\tilde{\pi}_{r}(\tilde{x})\|^{2}\tilde{\pi}_{r}(\tilde{x})\mathrm{d}\tilde{x}\right)^{1/2}$
[…]
where we use Cauchy’s inequality in the first step. This completes the proof. ∎
B.1 Lemma
Lemma B.1.
Under 4.1, it holds that (denote $\kappa=\frac{M}{m}$)
$\max_{j\in V}\sum_{k\in V}\|C^{1/2}(j,k)\|\leq m^{-1/2}S\nu!\kappa^{\nu}.$ (B.5)
Proof.
By Taylor expansion,
$x^{-1/2}=M^{-1/2}\left(1+\sum_{q=1}^{\infty}\frac{(2q-1)!!}{(2q)!!}\left(1-\frac{x}{M}\right)^{q}\right),\quad x\in[m,M].$
For $j\neq k$, let $n=\mathsf{d}_{G}(j,k)-1$, plug in $x=C^{-1}$ and notice $mI\preceq C^{-1}\preceq MI$, we obtain
$C^{1/2}=P_{n}(C^{-1})+M^{-1/2}\sum_{q=n+1}^{\infty}\frac{(2q-1)!!}{(2q)!!}\left(I-\frac{C^{-1}}{M}\right)^{q}.$
where $P_{n}\in\mathbb{P}_{n}$. By assumption, $C^{-1}(i,j)=0$ if $j\notin\mathcal{N}_{i}$. Since $\mathsf{d}_{G}(j,k)=n+1$, it holds
$P_{n}(C^{-1})(j,k)=0,$
as all the path connecting $j$ and $k$ has length no less than $\mathsf{d}_{G}(j,k)>n$. Therefore,
$\|C^{1/2}(j,k)\|=~{}$ $M^{-1/2}\left\|\sum_{q=n+1}^{\infty}\frac{(2q-1)!!}{(2q)!!}\left(I-\frac{C^{-1}}{M}\right)^{q}(j,k)\right\|$
[…]
For the coefficients, we bound
$\frac{(2q-1)!!}{(2q)!!}=~{}$ $\frac{1}{2}\prod_{j=2}^{q}\left(1-\frac{1}{2j}\right)\leq\frac{1}{2}\prod_{j=2}^{q}\exp\left(-\frac{1}{2j}\right)=\frac{1}{2}\exp\left(-\frac{1}{2}\sum_{j=2}^{q}\frac{1}{j}\right)$
[…]
Since $mI\preceq C^{-1}\preceq MI$, it holds that
$0\preceq\left(I-\frac{C^{-1}}{M}\right)^{q}\preceq\left(1-\frac{m}{M}\right)^{q}I~{}\Rightarrow~{}\left\|\left(I-\frac{C^{-1}}{M}\right)^{q}\right\|_{\rm op}\leq(1-\kappa^{-1})^{q}.$
For diagonal entry, $\|C^{1/2}(j,j)\|\leq\left\|C^{1/2}\right\|\leq m^{-1/2}$, so that
$\sum_{k\in V}\|C^{1/2}(j,k)\|=~{}$ $\|C^{1/2}(j,j)\|+\sum_{k\neq j}\|C^{1/2}(j,k)\|$
[…]
Here the second inequality follows the sparsity assumption, the third inequality follows A.2, and the last inequality follows $\Gamma\left(\nu+\frac{1}{2}\right)\leq\nu!$ and $s\nu!\kappa^{\nu-1}\geq 1$. This completes the proof. ∎
Appendix C Proof of 5.1
Proof.
By 2.1, $\pi$ is $\delta$-localized with $\delta=\frac{S\nu!\kappa^{\nu}}{m}$. 3.1 then implies
$\begin{split}\max_{i\in V}\mathsf{W}_{1}(p_{i},\hat{p}_{i})\leq~{}&\delta\cdot\max_{j\in V}\left\|\nabla_{j}\log p-\nabla_{j}\log\hat{p}\right\|_{L^{1}(p)}\\ \leq~{}&\delta\cdot\left(\max_{j\in V}\mathbb{E}_{X\sim p}\|s_{j}(X)-s_{\hat{\theta}_{*},j}(X)\|^{2}\right)^{1/2}=\delta\cdot J_{\rm loc}^{1/2}(\hat{\theta}_{*}).\end{split}$ (C.1)
Using optimality of $\hat{\theta}_{*}$ and $J_{\rm loc}(\theta_{*})=0$,
$\begin{split}J_{\rm loc}(\hat{\theta}_{*})=~{}&\left(J_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\hat{\theta}_{*})\right)+\left(\hat{J}_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\theta_{*})\right)+\left(\hat{J}_{\rm loc}(\theta_{*})-J_{\rm loc}(\theta_{*})\right)\\ \leq~{}&\left(J_{\rm loc}(\hat{\theta}_{*})-\hat{J}_{\rm loc}(\hat{\theta}_{*})\right)+0+\left(\hat{J}_{\rm loc}(\theta_{*})-J_{\rm loc}(\theta_{*})\right)\\ \leq~{}&2\max_{\theta\in\Theta}\|\hat{J}_{\rm loc}(\theta)-J_{\rm loc}(\theta)\|\\ =~{}&2\max_{\theta\in\Theta}\left|\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(\hat{J}_{j}(\theta)+\lambda_{j}\right)-\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(J_{j}(\theta)+\lambda_{j}\right)\right|\\ \leq~{}&2\max_{\theta\in\Theta}\max_{j\in V}|\hat{J}_{j}(\theta)-J_{j}(\theta)|.\end{split}$ (C.2)
Given data $X_{1},\dots,X_{N}$, denote for simplicity
$\mathbb{P}_{N}f=\frac{1}{N}\sum_{i=1}^{N}f(X_{i}),\quad f_{\theta,j}(x)=2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(x)\right)+\left\|s_{\theta,j}(x)\right\|^{2}.$
Then
$\hat{J}_{j}(\theta)-J_{j}(\theta)=(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta,j}.$
For any $\theta,\phi\in\Theta$, it holds that
$|f_{\theta,j}-f_{\phi,j}$ $|\leq 2|\Delta_{jj}\log p_{\theta}-\Delta_{jj}\log p_{\phi}|+|\left\|\nabla_{j}\log p_{\theta}\right\|^{2}-\left\|\nabla_{j}\log p_{\phi}\right\|^{2}|$
[…]
where we use $|\mathcal{N}_{j}|\leq S+1$ and $\left\|\nabla_{j}u_{\theta,k}(x)\right\|\leq R$.
Consider the function space
$\mathcal{U}_{\theta}:=\{u_{\theta}=(u_{\theta,j})_{j\in V}:u_{\theta,j}\in\mathcal{U}_{\theta,j}\}.$
Equip $\mathcal{U}_{\theta}$ with norm
$\left\|u_{\theta}\right\|_{\mathcal{U}}:=\max_{j\in V}\left\|u_{\theta,j}\right\|_{C^{2}}.$
By definition, $\sup_{u_{\theta}\in\mathcal{U}_{\theta}}\left\|u_{\theta}\right\|_{\mathcal{U}}\leq R$. Consider a $\varepsilon$-net $\mathcal{U}_{\varepsilon}=\{u_{\theta_{i}}\}_{n=1}^{N_{\varepsilon}}\subset\mathcal{U}_{\theta}$ of $\mathcal{U}_{\theta}$ with respect to $\left\|\cdot\right\|_{\mathcal{U}}$, i.e. $\forall u_{\theta}\in\mathcal{U}_{\theta}$, there exists some $u_{\theta_{i}}\in\mathcal{U}_{\varepsilon}$ s.t. $\left\|u_{\theta}-u_{\theta_{i}}\right\|_{\mathcal{U}}\leq\varepsilon$. Then
$\left\|f_{\theta,j}-f_{\theta_{i},j}\right\|_{L^{\infty}}\leq 2(S+1)R\sum_{k\in\mathcal{N}_{j}}\left\|u_{\theta,k}-u_{\theta_{i},k}\right\|_{C^{2}}\leq 2(S+1)^{2}R\varepsilon.$
Therefore,
$|\hat{J}_{j}(\theta)-J_{j}(\theta)|=~{}$ $|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta,j}|\leq|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|+|(\mathbb{P}_{N}-\mathbb{E}_{p})\left(f_{\theta,j}-f_{\theta_{i},j}\right)|$
[…]
This further implies that
$\max_{\theta\in\Theta}\max_{j\in V}|\hat{J}_{j}(\theta)-J_{j}(\theta)|\leq\max_{1\leq i\leq N_{\varepsilon}}\max_{j\in V}|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|+2(S+1)^{2}R\varepsilon.$
Next we give a bound on $N_{\varepsilon}$. It is a classical result (see for instance [43]) that the covering number of the unit ball $B_{1}(C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}})):=\{u\in C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}}):\left\|u\right\|_{C^{2}}\leq 1\}$ is bounded by
$\log N(\varepsilon,B_{1}(C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}})),L^{\infty})\leq C_{d_{\mathcal{N}_{j}}}\varepsilon^{-d_{\mathcal{N}_{j}}/2}.$
Here $\mathcal{U}_{\theta}$ is the product space of $C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}})$ equipped with the $\ell_{\infty}$-norm, so that
$\log N_{\varepsilon}\leq\sum_{j\in V}\log N(\varepsilon,B_{1}(C^{2}(\mathbb{I}^{d_{\mathcal{N}_{j}}})),L^{\infty})\leq bC_{d_{\rm loc}}\varepsilon^{-d_{\rm loc}/2}.$ (C.3)
Recall here $d_{\rm loc}=\max_{j}d_{\mathcal{N}_{j}}$. Notice $f_{\theta,j}$ is uniformly bounded by
$|f_{\theta,j}|\leq 2(S+1)R\sum_{k\in\mathcal{N}_{j}}\left\|u_{\theta,k}\right\|_{C^{2}}\leq 2(S+1)^{2}R^{2}.$
By Hoeffding’s inequality,
$\mathbb{P}\left[|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq\mathrm{e}^{-Nt^{2}/2(S+1)^{4}R^{4}},\quad\forall t>0.$
So that by maximal inequality,
$\mathbb{P}\left[\max_{1\leq i\leq N_{\varepsilon}}\max_{j\in V}|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq N_{\varepsilon}b\cdot\mathrm{e}^{-Nt^{2}/2(S+1)^{4}R^{4}}.$
Take $t=(S+1)^{2}R^{2}\varepsilon$, then
$\mathbb{P}\left[\max_{\theta\in\Theta}\max_{j\in V}|\hat{J}_{j}(\theta)-J_{j}(\theta)|\geq 3t\right]\leq\mathbb{P}\left[\max_{1\leq i\leq N_{\varepsilon}}\max_{j\in V}|(\mathbb{P}_{N}-\mathbb{E}_{p})f_{\theta_{i},j}|\geq t\right]\leq N_{\varepsilon}b\cdot\mathrm{e}^{-N\varepsilon^{2}/2}.$
Recall (C.2) and (C.3), we obtain
$\mathbb{P}\left[J_{\rm loc}(\hat{\theta}_{*})\geq 6(S+1)^{2}R^{2}\varepsilon\right]\leq N_{\varepsilon}b\cdot\mathrm{e}^{-N\varepsilon^{2}/2}\leq b^{2}C_{d_{\rm loc}}\varepsilon^{-d_{\rm loc}/2}\mathrm{e}^{-N\varepsilon^{2}/2}.$
In order to make the right hand side less than $\lambda\in(0,1)$, one can take
$\varepsilon^{2}=\frac{C}{N}\left(\log\lambda^{-1}+\log b\right),$
where $C$ is some constant depending only on $d_{\rm loc}$. So that with probability at least $1-\lambda$,
$J_{\rm loc}(\hat{\theta}_{*})\leq 6(S+1)^{2}R^{2}\varepsilon\leq C^{\prime}N^{-1/2}\sqrt{\log\lambda^{-1}+\log b}.$
The conclusion follows by plugging in the bound (C.1) on $\max_{j}\mathsf{W}_{1}(p_{j},\hat{p}_{j})$. ∎